### v3

# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks but remain fundamentally static, unable to adapt their internal parameters to novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this static nature has become a critical bottleneck, necessitating agents that can adaptively reason, act, and evolve in real time. This paradigm shift —from scaling static models to developing self-evolving agents — has sparked growing interest in architectures and methods enabling continual learning and adaptation from data, interactions, and experiences. This survey provides the first systematic and comprehensive review of self-evolving agents, organizing the field around three foundational dimensions — what to evolve, when to evolve, and how to evolve. We examine evolutionary mechanisms across agent components (e.g., models, memory, tools, architecture), categorize adaptation methods by stages (e.g., intra-test-time, inter-test-time), and analyze the algorithmic and architectural designs that guide evolutionary adaptation (e.g., scalar rewards, textual feedback, single-agent and multi-agent systems). Additionally, we analyze evaluation metrics and benchmarks tailored for self-evolving agents, highlight applications in domains such as coding, education, and healthcare, and identify critical challenges and research directions in safety, scalability, and co-evolutionary dynamics. By providing a structured framework for understanding and designing self-evolving agents, this survey establishes a roadmap for advancing adaptive, robust, and versatile agentic systems in both research and real-world deployments, ultimately shedding lights to pave the way for the realization of Artificial Super Intelligence (ASI), where agents evolve autonomously, performing at or beyond human-level intelligence across a wide array of tasks.
Figure 1: A conceptual trajectory illustrating the progression from large language models (LLMs) to foundation agents, advancing to self-evolving agents—our focus, and ultimately toward hypothetical Artificial Super Intelligence (ASI). Along this path, intelligence and adaptivity increase, marking a shift toward more autonomous and agentic AI systems.
1 Introduction
"It is not the most intellectual of the species that survives; it is not the strongest that survives; but the species that survives is the one that is able best to adapt and adjust to the changing environment in which it finds itself."                          — Charles Darwin [^1] This quote is widely attributed to Charles Darwin, but it does not appear verbatim in his writings. The phrasing is believed to originate from Professor Leon C. Megginson, who paraphrased Darwin’s ideas. Despite its frequent misattribution, the quote effectively captures the essence of Darwinian evolution and has since been popularized in both scientific and managerial literature.
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Yet, they remain fundamentally static luo2025largelanguagemodelagent , unable to adapt their internal parameters when encountering novel tasks, evolving knowledge domains, or dynamic interaction contexts. As LLMs are increasingly deployed in open-ended, interactive environments, this limitation becomes a critical bottleneck. In such settings, conventional knowledge retrieval mechanisms prove inadequate, giving rise to agents capable of dynamically adapting their perception, reasoning, and actions in real time. This emerging need for dynamic, continual adaptation signals a conceptual shift in artificial intelligence: from scaling up static models to developing self-evolving agents, such agents are capable to continuously learn from new data, interactions, and experiences in real-time, leading to systems that are more robust, versatile, and capable of tackling complex, dynamic real-world problems (tool_tut, ). This shift is currently driving us toward a promising and transformative path to Artificial Super Intelligence (ASI), where the agents not only can learn and evolve from experience with an unpredictable speed but also perform at or above human-level intelligence across a wide array of tasks (wang2025theoryagentstoolusedecisionmakers, ).
Unlike static LLMs, which remain constrained by their inability to adapt to novel and evolving contexts, self-evolving agents are designed to overcome these limitations by continuously learning from real-world feedback. This progression reshapes our understanding of agents. Self-evolving agents, as a core concept, will be the precursors to ASI, acting as intermediaries that pave the way for the ultimate evolution of intelligence, as shown in Figure 1. Recent research initiatives have increasingly focused on developing adaptive agent architectures capable of continually learning and adapting from experience, such as recent advancements in agent frameworks (yin2025godelagentselfreferentialagent, ), prompting strategies fernando2023promptbreeder , and different optimization ways to evolve. Notwithstanding these advances, existing surveys predominantly address agent evolution as a subsidiary component within comprehensive agent taxonomies. Previous surveys primarily provide systematic overviews of general agent development, while offering limited coverage of self-evolving mechanisms across constrained scenarios in self-evolving agents (luo2025largelanguagemodelagent, ; liu2025advanceschallengesfoundationagents, ). For example, Luo et al. (luo2025largelanguagemodelagent, ) discuss several ways to evolve, such as self-learning and multi-agent co-evolution, while Liu et al. (liu2025advanceschallengesfoundationagents, ) explicitly introduce the evolution in terms of different components of agents, such as tools and prompts. Moreover, some studies focus specifically on the evolution of language models themselves tao2024survey-self-evo-llm , rather than on the broader concept of agents. Yet, there is no systematic survey devoted to a dedicated, comprehensive investigation of self-evolving agents as a first-class research paradigm. This gap has left fundamental questions underexplored: What aspects of an agent should evolve? When should adaptation occur? And how should that evolution be implemented in practice?
To the best of our knowledge, this is the first systematic and comprehensive survey focusing on self-evolving agents, offering a clear roadmap for both theoretical inquiry and practical deployment. We organize our analysis around three foundational questions — what, when, and how to evolve — and provide a structured framework for understanding each. Specifically, we systematically examine individual agent components, including the model, memory, tools and corresponding workflow, investigating their distinct evolutionary mechanisms (what to evolve of agent in Section 3); then we divide existing evolving methods according to different temporal stages with different learning paradigms such as supervised fine-tuning, reinforcement learning and inference-time evolving (when to evolve in Section 4). We finally summarize different signals to guide the evolution of agents, such as textual feedback or scalar rewards, and also different architectures of agents to evolve, such as single-agent and multi-agent evolution (how to evolve in Section 5). Furthermore, we review certain evaluation metrics and benchmarks to track existing advancements of self-evolving agents, emphasizing the importance of co-evolution between evaluation and agents (Section 6). We also examine emerging applications in domains such as coding, education, and healthcare, where continual adaptation and evolution are essential (Section 7). Finally, we identify persistent challenges and outline promising research directions to guide the development of self-evolving agents (Section 8). Through this systematic decomposition of self-evolutionary processes across orthogonal dimensions, we provide a structured and practical framework enabling researchers to systematically analyze, compare, and design more robust and adaptive agentic systems. To sum up, our key contributions are as follows:
  • We establish a unified theoretical framework for characterizing self-evolutionary processes in agent systems, anchored around three fundamental dimensions: what evolves, how it evolves, and when it evolves, providing clear design guidance for future self-evolving agentic systems.
  • We further investigate the evaluation benchmark or environment tailored for self-evolving agents, highlighting emerging metrics and challenges related to adaptability, robustness, and real-world complexity.
  • We showcase several key real-world applications across various domains, including autonomous software engineering, personalized education, healthcare, and intelligent virtual assistance, illustrating the practical potential of self-evolving agents.
  • We identify critical open challenges and promising future research directions, emphasizing aspects like safety, personalization, multi-agent co-evolution, and scalability.
In doing so, our survey provides researchers and practitioners with a more structured taxonomy for understanding, comparing, and advancing research of self-evolving agents from different perspectives. As LLM-based agents are increasingly integrated into mission-critical applications, understanding their evolutionary dynamics becomes essential, extending beyond academic research to encompass industrial applications, regulatory considerations, and broader societal implications.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=0.1em, edge+=darkgray, line width=0.8pt, s sep=2pt, inner xsep=2pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=3em,font=, where level=2text width=5em,font=, where level=3text width=3em,font=, where level=4text width=14em,font=, [ Self-evolving Agent, draw=gray, color=gray!100, fill=gray!15, thick, text=black, ver, [What to evolve, cyan!100, fill=cyan!8,thick, text=black, [Model [ Policy [ SCA zhou2025selfchallenginglanguagemodelagents , Self-Rewarding Self-Improving simonds2025selfrewardingselfimproving , SELF lu2023self , SCoRe kumar2024training , PAG jiang2025pag , TextGrad yellamraju2024textgrad , AutoRule wang2025autorule , SRLM wang2025selfreasoninglanguagemodelsunfold ]] [ Lesson [ SCA zhou2025selfchallenginglanguagemodelagents , AgentGen hu2024agentgen , Reflexion shinn2023reflexion , AdaPlanner sun2023adaplanner , SICA robeyns2025selfimprovingcodingagent , Self-Refine madaan2023selfrefine , Learn-by-interact su2025learnbyinteractdatacentricframeworkselfadaptive , RAGEN wang2025ragen , DYSTIL wang2025dystildynamicstrategyinduction ]]] [Context [ Memory [ SAGE liang2024sage , Mem0 chhikara2025mem0 , MemInsight salama2025meminsight , REMEMBER zhang2024large , Expel zhao2024expel , Agent Workflow Memory wang2024agent , Richelieu diplomacy agent guan2024richelieu , ICE qian2024investigate ]][ Prompt [ APE zhou2022large , ORPO yang2023large , ProTeGi pryzant2023automatic , PromptAgent wang2023promptagent , REVOLVE zhang2024revolve , PromptBreeder fernando2023promptbreeder , DSPy khattab2023dspy , Trace wang2023trace , TextGrad yellamraju2024textgrad , SPO xiang2025self , LLM-AutoDiff yin2025llm , EvoAgent yuan2024evoagent ]]] [Tool[ Creation [ Voyager wang2023voyager , Alita qiu2025alita , ATLASS Haque2025ATLASS , CREATOR qian2023creator , SkillWeaver zheng2025skillweaver , CRAFT yuan2024craft ]][ Mastery [ LearnAct zhao2024learnact , DRAFT fromexplorationtomastery , ToolLLM qin2023toolllm , Toolformer schick2023toolformer , Gorilla patil2024gorilla ]] [ Selection [ ToolGen wang2025toolgen , AgentSquare shang2025agentsquare , Darwin Godel Machine zhang2025darwin , COLT qu2024colt , TOOLRET shi2025toolret , ToolRerank zheng2024toolrerank , PTR gao2024ptr , SSO Nottingham2024sso ]]] [Architecture[ Single-Agent [ AgentSquare shang2025agentsquare , Darwin Godel Machine zhang2025darwin , Gödel Agent  yin2025godelagentselfreferentialagent , AlphaEvolve novikov2025alphaevolve , TextGrad yellamraju2024textgrad , EvoFlow zhang2025evoflow , MASS zhou2025maasprompt ]][ Multi-Agent [ AFlow zhang2024aflow , ADAS hu2024automated , AutoFlow li2024autoflow , GPTSwarm zhuge2024gptswarm , ScoreFlow wang2025scoreflow , FlowReasoner gao2025flowreasoner , ReMA wan2025remalearningmetathinkllms , GIGPO feng2025groupingrouppolicyoptimizationllm ] ] ] ] [ When to evolve, color=green!80!black!50!, fill=green!5, thick, text=black, [ Intra-test-time self-evolution [ ICL [ Reflexion shinn2023reflexion , SELF madaan2023self_refine , AdaPlanner sun2023adaplanner , TrustAgent hua2024trustagent ] ] [ SFT [ Self-Adaptive LM zweiger2025self , TTT-NN hardt2023test , SIFT hubotter2024efficiently ] ] [ RL [ LADDER simonds2025ladder , Ttrl zuo2025ttrl ] ] ] [ Inter-test-time self-evolution [ SFT [ SELF lu2023self , STaR zelikman2022star , Quiet-STaR zelikman2024quiet , SiriuS zhao2025sirius , Chen et al. chen2025sft ] ] [ RL [ RAGEN wang2025ragen , Learning-Like-Humans zhang2025learning , WebRL qi2024webrl , DigiRL bai2024digirl ] ] ] ] [ How to evolve, color=yellow!80!brown, fill=yellow!20!white, thick, text=black [ Reward-based Self-Evolution [Textual Feedback [ Reflexion shinn2023reflexion , AdaPlanner sun2023adaplanner , AgentS2 agashe2025agents2 , SELF lu2023self , Self-Refine madaan2023self_refine , SCoRe kumar2024training , PAG jiang2025pag , TextGrad yellamraju2024textgrad ] ] [Internal Rewards [ CISC taubenfeld2025cisc , Self-Ensemble xu2025selfensemblemitigatingconfidencedistortion , SRSI simonds2025selfrewardingselfimproving , Self-Certainty kang2025scalablebestofnselectionlarge , Self-Rewarding Language Models yuan2025selfrewardinglanguagemodels ] ] [External Rewards [ Self-Train LM shafayat2025largereasoningmodelsselftrain , MM-UPT wei2025unsupervisedposttrainingmultimodalllm , CoVo zhang2025consistentpathsleadtruth , SWE-Dev du2025swedevevaluatingtrainingautonomous , SICA robeyns2025self , Feedback Friction jiang2025feedbackfriction , USEagent applis2025useagent , DYSTIL wang2025dystildynamicstrategyinduction , OTCPO wang2025otcpo , AutoRule wang2025autorule , EGSR zhang2025learning , LADDER simonds2025ladder , RAGEN wang2025ragen , SPIRAL liu2025spiral ] ] [Implicit Rewards [ Reward Is Enough song2025reward , Endogenous reward li2025generalist ] ] ] [ Imitation & Demonstration Learning [Self-Generated [ STaR zelikman2022star , V-STaR hosseini2024v , AdaSTaR koh2025adastar , STIC deng2024enhancing , GENIXER zhao2024genixer ] ] [Others [ SiriuS zhao2025sirius , SOFT tang2025bridging , RISE qu2024recursive , IoE li2024confidence ] ] ] [ Population-based & Evolutionary Methods [Single Agent [ DGM zhang2025darwin , GENOME zhang2025nature , SPIN chen2024self , SPC chen2025spc , STL mendes2025language ] ] [Multi-Agent [ EvoMAC hu2024self , Puppeteer dang2025multi , MDTeamGPT chen2025self , MedAgentSim almansoori2025self ] ] ] ] [ Where to Evolve, color=violet!60, fill=violet!15, thick, text=black, [ General Domain [ Memory [ Mobile-Agent-E wang2025mobile , MobileSteward liu2025mobilesteward ] ] [ Curriculum Learning [ WebRL qi2024webrl , Voyager wang2023voyager ] ] [ Model-Agent Co-Evolution [ WebEvolver fang2025webevolver , UI-Genie xiao2025ui , Absolute-Zero zhao2025absolute ] ] ] [ Specialized Domain [ Coding, GUI Finalicial… [ EvoMac hu2024self , SICA robeyns2025self , AgentCoder dang2025multi , QuantAgent wang2024quantagent ] ] [ Others [ Arxiv-copilot lin2024arxivcopilot , Voyager wang2023voyager ] ] ] ] ]
Figure 2: Taxonomy of self-evolving agents, in which agents are analyzed along the what, when, how, and where dimensions, with selected representative methods and systems annotated at each leaf node.
Figure 3: A comprehensive overview of self-evolving agents across key dimensions: What to evolve: covering four major categories—Model, Context, Tool, and Architecture; When to evolve: differentiating intra-test-time and inter-test-time self-evolution, via in-context learning (ICL), supervised fine-tuning (SFT), or reinforcement learning (RL); How to evolve: centered on three main paradigms—reward-based, imitation and demonstration, and population-based methods. These are complemented by cross-cutting dimensions. Where to evolve: ranging from general-purpose domains to specific domains; Evaluation: focusing on goals (e.g., adaptivity, safety, generalization) and evaluation paradigms (static, short-horizon, or long-horizon).
Figure 4: An evolutionary landscape of several representative self-evolving agent frameworks from 2022 to 2025. The figure chronologically organizes major research milestones in the development of self-evolving agents with capabilities such as autonomous planning, tool use, and continual self-improvement.
2 Definitions and Foundations
Before delving into a comprehensive survey, we first present a formal definition of self-evolving agents and introduce a taxonomy of the key aspects in self-evolving agents. We also discuss the relationships between self-evolving agents and other renowned learning paradigms, such as curriculum learning, lifelong learning, model editing, and unlearning, highlighting the adaptive, dynamic, and autonomous nature of self-evolving agents.
2.1 Definitions
Environment
We first define the environment (including the user and the execution environment, e.g., Linux shell) of an agent system as a partially observable Markov Decision Process (POMDP), represented as a tuple $E=(\mathcal{G},\mathcal{S},\mathcal{A},T,R,\Omega,O,\gamma)$, where:
  • $\mathcal{G}$ is a set of potential goals. Each $g\in\mathcal{G}$ is a task objective that the agent needs to achieve, e.g., a user query.
  • $\mathcal{S}$ is a set of states. Each $s\in\mathcal{S}$ represents the internal state of the environment.
  • $\mathcal{A}$ is a set of actions. Each action $a\in\mathcal{A}$ can be a combination of textual reasoning, retrieval of external knowledge, and tool calls.
[…]
  • $R:\mathcal{S}\times\mathcal{A}\times\mathcal{G}\rightarrow\mathcal{R}$ is the feedback/reward function, conditioned on the specific goal $g\in\mathcal{G}$. The feedback $r=R(s,a,g)$ typically takes the form of a scalar score or textual feedback.
  • $\Omega$ is a set of observations accessible to the agent.
[…]
  • $\gamma$ is the discount factor.
Agent system
We define a (multi-)agent system as $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$. The architecture $\Gamma$ determines the control flow of the agent system or collaborative structures between multiple agents. It is typically represented as a sequence of nodes $(N_{1},N_{2},...)$ organized by graph or code structures. Each node $N_{i}$ consists of the following components:
  • $\psi_{i}$: the underlying LLM/MLLM.
  • $C_{i}$: the context information, e.g., prompt $P_{i}$ and memory $M_{i}$.
  • $\mathcal{W}_{i}$: the set of available tools/APIs.
At each node, the agent policy is a function $\pi_{\theta_{i}}(\cdot|o)$ that takes an observation and outputs the probability distribution of the next action, where $\theta_{i}=(\psi_{i},C_{i})$. The actual action space here is the union of the natural language space and the tool space $\mathcal{W}_{i}$.
For a given task $\mathcal{T}=(E,g)$, represented by an environment $E$ and a corresponding goal $g\in\mathcal{G}$, the agent system follows the topology $\Gamma$ to generate a trajectory $\tau=(o_{0},a_{0},o_{1},a_{1},...)$, and receives a feedback $r$ either from the external environment or from internal signals (e.g., self-confidence or feedback from an evaluator).
Self-evolving strategy
A self-evolving strategy is a transformation $f$ that maps the current agent system to a new state, conditioned on the generated trajectory $\tau$ and the external/internal feedback $r$:
$f(\Pi,\tau,r)=\Pi^{\prime}=(\Gamma^{\prime},\{\psi^{\prime}_{i}\},\{C^{\prime}_{i}\},\{\mathcal{W}^{\prime}_{i}\})$ (1)
Objective of self-evolving agents
Let $U$ be a utility function that measures the performance of an agent system $\Pi$ on a given task $\mathcal{T}$ by assigning a scalar score $U(\Pi,\mathcal{T})\in\mathbb{R}$. The utility may be derived from the task-specific feedback $r$, such as a reward signal or textual evaluation, possibly combined with other performance indicators (e.g., completion time, accuracy, or robustness). Given a sequence of tasks $(\mathcal{T}_{0},\mathcal{T}_{1},...,\mathcal{T}_{n})$ and an initial agent system $\Pi_{0}$, a self-evolving strategy $f$ recurrently generates an evolving sequence of agent systems $(\Pi_{1},\Pi_{2},...,\Pi_{n})$ via
$\Pi_{j+1}=f(\Pi_{j},\tau_{j},r_{j}),$ (2)
where $\tau_{j}$ and $r_{j}$ are the trajectory and feedback on task $\mathcal{T}_{j}$.
The overarching objective in designing a self-evolving agent is to construct a strategy $f$ such that the cumulative utility over tasks is maximized:
$\max_{f}\sum_{j=0}^{n}U(\Pi_{j},\mathcal{T}_{j})$ (3)
2.2 Relationships with Other Works
Table 1 summarizes the key distinctions between self-evolving agents and other paradigms (including curriculum learning, lifelong learning, model editing, and unlearning). Unlike these existing paradigms that primarily focus on updating model parameters, self-evolving agents expand the scope of updating targets to include non-parametric components, such as context (prompts and memory) and toolset. This expanded space provides greater flexibility, enabling self-evolving agents to operate effectively in sequential task settings and adapt at test time. More crucially, self-evolving agents uniquely demonstrate the ability of active exploration (e.g., searching for open-source tools online qiu2025alita ), structural modification of their own topology (e.g., iteratively modifying the workflow zhang2024aflow or code zhang2025darwin ), and self-reflection and self-evaluation capabilities (e.g., providing verbal feedback using an internal evaluator LLM shinn2023reflexion ), which are absent in previous paradigms.
We provide a brief introduction to each paradigm below, highlighting the differences among these paradigms, as well as the differences with self-evolving agents.
Curriculum Learning
Curriculum learning is a training strategy for AI models in which data are presented in order of increasing difficulty bengio2009curriculum ; wang2021survey . This strategy resembles human curricula where concepts are introduced progressively from simple to complex. Curriculum learning has been widely adopted across diverse domains, including computer vision guo2018curriculumnet ; jiang2014easy ; liu2023towards , natural language processing platanios2019competence ; tay2019simple , speech recognition braun2017curriculum ; lotfian2019curriculum , etc. Recently, several curriculum learning-based methods have been proposed to fine-tune LLMs during the post-training phase wang2025dump ; zhang2025preference ; parashar2025curriculum ; zhang2025learning ; li2025curriculum . The framework for curriculum learning generally comprises two key components: a difficulty measurer that quantifies the difficulty level of each training data point, and a training scheduler that reorganizes the order of data points received by the model according to the difficulty level. Unlike curriculum learning, which operates on a static dataset, self-evolving agents aim to handle sequential tasks in dynamic environments. Additionally, curriculum learning updates only model parameters, whereas self-evolving agents are able to adjust non-parametric components like memory and tools.
Lifelong Learning
Lifelong learning refers to the ability of AI models to continuously and adaptively learn when exposed to new tasks and environments, while retaining previously acquired knowledge and abilities. This learning paradigm, also known as continual learning or incremental learning, is crucial for AI models to operate in dynamic and complex environments wang2024comprehensive ; zheng2025lifelong ; parisi2019continual ; shi2024continual ; yang2025recent ; zhou2024continual . The primary goal of lifelong learning for AI models is to achieve a balance between preserving existing knowledge (stability) and acquiring new knowledge (plasticity) when exposed to new data or tasks mccloskey1989catastrophic ; zheng2025lifelong ; ratcliff1990connectionist ; rolnick2019experience . Though it shares the sequential task setting with self-evolving agents, lifelong learning differs in two ways: (1) like curriculum learning, lifelong learning typically updates only model parameters, lacking the ability to modify non-parametric components; (2) lifelong learning primarily acquires knowledge passively through external feedback or manual guidance, whereas self-evolving agents actively explore their environment and may incorporate internal reflection or self-evaluation mechanisms.
Model Editing and Unlearning
Model editing and unlearning aim to efficiently and precisely modify specific knowledge in AI models while preserving irrelevant knowledge and avoiding full retraining wang2024knowledge ; wang2025comprehensive ; zhang2024comprehensive ; wang2025comprehensive ; nguyen2022survey ; geng2025comprehensive . A canonical application of model editing is to perform efficient and precise localized factual updates (e.g., modifying the answer to “2021 Olympics host city” from "Tokyo" to "Paris"). Early methods focused on triples of atomic knowledge and later expanded into various trustworthy-related tasks fang2025alphaedit ; huang2025model . Recent studies also propose lifelong model editing chen2024lifelong that sequentially performs model editing. For model unlearning, early efforts mainly focus on the removal of privacy-related information chen2021machine . With the rapid development of LLMs, model unlearning is also used to enhance LLMs’ safety safe_unlearning ; li2024wmdp ; circuit_breaker ; lu2025x . Compared to lifelong learning, model editing shares an aligned objective: both aim to acquire new knowledge or capabilities while mitigating catastrophic forgetting. However, lifelong learning typically relies on extensive gradient-based fine-tuning across all model parameters, whereas model editing often modifies only a small subset of parameters in a targeted manner. Compared to self-evolving agents, model editing (1) cannot modify non-parametric components such as memory or tools, and (2) relies on a pre-defined pipeline from the algorithm designer, whereas self-evolving agents can spontaneously employ more diverse and flexible strategies based on the observation of the environment or internal feedback signals.
Table 1: Comparison between self-evolving agents and other renowned paradigms, including curriculum learning, lifelong learning, model editing and unlearning.
Paradigm Evolving Evolving Dynamic Test-time Active Structural Self-reflect
Context Toolset Tasks Adaptation Exploration Change & Eval
Curriculum Learning
Lifelong Learning
Model Editing
Self-evolving Agents
3 What to Evolve?
The transition from pre-programmed, static systems to genuinely autonomous intelligent agents hinges on one critical capability: self-evolution. This capability for continuous improvement is not monolithic; instead, it manifests across various facets of an agent’s being. Therefore, the first key aspect of a self-evolving agent is identifying the evolving components – which components in the agent system $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$ can be explicitly changed over time?
Following the formulation in Section 2.1, we can decompose the agent system into four fundamental, evolvable pillars. Our investigation starts at the agent’s cognitive core, namely the Models $\{\psi_{i}\}$, and examines how the fundamental parameters that govern its reasoning and behavior are continuously updated through its own experienceszhou2025selfchallenginglanguagemodelagents ; wang2025ragen . We then consider the Context $\{C_{i}\}$ that shapes its actions, exploring the evolution of both the instructionsxiang2025self ; khattab2023dspy it follows and the long-term memorychhikara2025mem0 ; wang2024agent it draws upon to stay informed and adapt. From this internal foundation, we shift to the agent’s external capabilities, analyzing how it evolves its set of Tools $\{\mathcal{W}_{i}\}$ by autonomously creatingqiu2025alita , masteringfromexplorationtomastery , and managing new skillswang2025toolgen to overcome its innate limitations. Finally, we scale our perspective to the Agentic System itself, investigating how the agentic system’s architecturehu2024automated ; zhang2024aflow and collaborative structureswan2025remalearningmetathinkllms are dynamically optimized to enhance overall performance and efficiency. We present a subset of these evolving dimensions in Table 2.
3.1 Models
Models constitute the central substrate of intelligent agents, directly determining their reasoning, planning, and decision-making behaviors. The ability of these models to evolve by continually adapting their internal parameters and expanding their functional capabilities is essential for the development of autonomous, general-purpose agents. Unlike static systems that rely heavily on human-annotated datasets and fixed training regimes, self-evolving models can improve through interaction, self-supervised data generation, and dynamic learning loops, thereby achieving greater efficiency, adaptability, and scalability. In detail, we outline the principal axes along which model evolution unfolds. These include learning from self-generated supervision to refine model weights, evolving through interaction with constructed or external environments. Together, these strategies represent a shift from passive learning paradigms toward active, continual, and self-directed improvement.
Table 2: Representative self-evolving agent methods positioned along four evolutionary pillars; a filled bullet ($\bullet$) marks dimensions where the approach actively evolves.
Method Model Context Tool Architecture
Policy Experience Prompt Memory Creation Mastery Selection Single Multi
SCA zhou2025selfchallenginglanguagemodelagents $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$
RAGEN wang2025ragen $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AgentGen  hu2024agentgen $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$
Promptbreeder fernando2023promptbreeder $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
Expel zhao2024expel $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Agent Workflow Memory wang2024agent $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$
Mem0 chhikara2025mem0 $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
MAS-Zero ke2025maszero $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
Multi-Agent Design zhou2025maasprompt $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$
SPO xiang2025self $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Alita qiu2025alita $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
TextGrad yellamraju2024textgrad $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$
DGM zhang2025darwin $\circ$ $\circ$ $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
AlphaEvolve novikov2025alphaevolve $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$
ADAS hu2024automated $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\bullet$
AFlow zhang2024aflow $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\bullet$
ReMA wan2025remalearningmetathinkllms $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$
SkillWeaver zheng2025skillweaver $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
LearnAct zhao2024learnact $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\circ$
DRAFT fromexplorationtomastery $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
ToolGen wang2025toolgen $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
CRAFT yuan2024craft $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\circ$
CREATOR qian2023creator $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$ $\circ$ $\circ$ $\circ$
Voyager wang2023voyager $\circ$ $\circ$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\circ$ $\bullet$
Policy
A self-evolving agent can refine its parameters to perform better on targeted tasks. Traditional methods of data collection for training agents on tool-use benchmarks are costly and often yield limited coverage, while purely synthetic data-generation pipelines typically suffer from inadequate quality. Consequently, recent studies emphasize enabling agents to autonomously generate data to improve their own model weights. One representative approach is the Self-Challenging Agent (SCA)zhou2025selfchallenginglanguagemodelagents , where a language model alternates roles between a challenger generating executable Code-as-Task problems and an executor solving them. The model then fine-tunes its parameters using trajectories derived from successful solutions, resulting in significant performance gains on complex, multi-step tasks. Similarly, the Self-Rewarding Self-Improving frameworksimonds2025selfrewardingselfimproving implements an internal self-judging mechanism, allowing the model to autonomously generate problems, solve them, and assess its performance, thus producing self-contained fine-tuning data without external annotations. This method demonstrated notable improvements, particularly in complex reasoning tasks. Beyond task creation, another promising research direction involves leveraging interaction feedback directly for parameter updates. For instance, SELF lu2023self , SCoRe kumar2024training , and PAG jiang2025pag interpret execution traces or natural-language critiques as reward signals within an online Supervised Fine-Tuning (SFT) combined with Reinforcement Learning (RL) framework, enabling continuous policy improvement. TextGrad yellamraju2024textgrad further extends this concept by treating unstructured textual feedback as a differentiable training signal capable of directly influencing both prompt design and model parameters. Additionally, AutoRule wang2025autorule converts language-model reasoning traces and preference feedback into explicit rule-based training rewards, enhancing the quality of model outputs through structured reward signals. Collectively, these advancements chart a clear trajectory—from agents autonomously crafting their training tasks to directly refining their parameters based on execution feedback, highlighting the capacity of models to evolve continuously by learning from the data they produce.
Experience
Agents can evolve not only by adjusting their internal parameters but also by actively interacting with or even constructing their environments, capturing experiences, and transforming them into learning signals that drive iterative improvement. This environmental loop provides agents with the complexity and diversity required for scalable self-adaptation. The Self-Challenging Agent (SCA)zhou2025selfchallenginglanguagemodelagents exemplifies this dynamic at the task level, where the agent autonomously generates novel Code-as-Task problems, executes them, and then filters successful trajectories for retraining itself. AgentGenhu2024agentgen extends this concept to full-environment generation, synthesizing diverse simulation worlds (in PDDL or Gym-style formats) derived from an initial corpus. It implements a bidirectional evolution loop that progressively adjusts task difficulty, enabling the agent to continuously grow within a dynamically structured curriculum. Reflexion shinn2023reflexion complements this by introducing self-reflective mechanisms, where agents iteratively record natural-language critiques of their previous actions, guiding future behavior to avoid recurring mistakes. Additionally, AdaPlanner sun2023adaplanner introduces closed-loop adaptive planning, allowing agents to refine their strategies on-the-fly based on environmental feedback, effectively reshaping action sequences in response to immediate outcomes. Similarly, Self-Refine madaan2023selfrefine employs an iterative refinement loop in which the agent repeatedly critiques and revises its initial outputs, significantly improving task accuracy without explicit retraining. SICA (Self-Improving Coding Agent) robeyns2025selfimprovingcodingagent further pushes the boundary by enabling agents to autonomously edit their underlying code and tools, iteratively enhancing their core reasoning abilities through direct self-modification. From a reinforcement learning perspective, frameworks such as RAGEN wang2025ragen and DYSTIL wang2025dystildynamicstrategyinduction conceptualize multi-step tool-use tasks as Markov Decision Processes, optimizing agent policies through rich environmental rewards and strategy induction loops. RAGEN leverages dense feedback from the environment to iteratively fine-tune action policies, while DYSTIL utilizes high-level strategy advice generated by language models to progressively internalize complex decision-making skills into reinforcement learning agents. Collectively, these approaches highlight a compelling paradigm where self-evolving agents not only leverage self-generated data but actively reshape their environments and internal mechanisms to fuel ongoing learning. Such dynamic interaction loops point toward autonomous, open-ended improvement cycles deeply grounded in experiential adaptation.
3.2 Context
An essential component of an LLM agent to be evolved is the context, which shapes how an agent behaves. To start with, we want to interpret two terms, "prompt optimization" and "memory evolution", which have been used in different literature. In most cases, these two terms can be used interchangeably because they both refer to what is included in the context window. Prompt optimization asks "how can we phrase or structure the instructions so the LLM behaves better?", and attends to details such as the wording, ordering. On the other hand, memory evolution asks "how should we store, forget, and retrieve context so that the agent can stay informed and perform better?", which focuses on what past information to surface or archive.
3.2.1 Memory Evolution
LLM-based agents are increasingly designed with long-term memory mechanisms that grow and adapt as the agent continues to solve tasks and interacts with its environment shan2025cognitive ; qian2023experiential . An evolving memory enables the agent to accumulate knowledge, recall past events, and adjust its behavior based on experience. Many works stress that effective memory management is crucial for agent performance zhong2024memorybank ; zhang2025g ; yan2024efficient . SAGE liang2024sage uses the Ebbinghaus forgetting curve to decide what to remember or forget. A-memxu2025mem updates the agent memory structure to create interconnected knowledge networks through dynamic indexing and linking, following the basic principles of the Zettelkasten method. Mem0 chhikara2025mem0 introduces a two-phase pipeline where the agent first extracts salient facts from recent dialogue and then decides how to update the long-term memory: the agent can ADD new facts, MERGE/UPDATE redundant ones, or DELETE contradictions. Such a mechanism ensures the agent’s long-term memory is coherent and up-to-date. MemInsight salama2025meminsight augments raw memories with semantic structure, which summarizes and tags past interactions for retrieval later. REMEMBER zhang2024large combines an LLM with a memory of experiences and uses reinforcement learning signals to decide how to update that memory after each episode.
A critical aspect of memory evolution is enabling agents to learn heuristics or skills from past experiences. Rather than only retrieving exact past instances, advanced agents distill experiences into more general guidance zhao2024expel ; fu2024autoguide . Expel zhao2024expel processes past trajectories to generate insights and rules to guide further interactions. This experiential knowledge accumulation leads to measurable gains, as the agent steadily performs better with more experience. Other systems focus on storing higher-level building blocks of problem-solving. For instance, Agent Workflow Memory wang2024agent records common sub-task sequences (workflows) so that an agent solving a complex task can retrieve and reuse a proven sequence of actions rather than plan from scratch. In the Richelieu diplomacy agent, the system improves its negotiation strategies by augmenting its memory through self-play games, storing the insights from simulated interactions to refine future decisions guan2024richelieu . By generalizing from specific episodes to reusable knowledge, these approaches illustrate how memory evolution turns an agent’s one-time experiences into long-term competencies, which leads to agents evolving.
3.2.2 Prompt Optimization
While memory evolution focused on what knowledge an agent retains, Prompt Optimization (PO) enables LLM agents to self-evolve by refining the instructions it feeds to the backbone model, which directly alters the model’s behavior without modifying model weights ramnath2025systematic . Early research treats instruction design as a search problem. APE zhou2022large generates candidate prompts, scores them on validation examples, and selects the best. ORPO yang2023large extends this idea by letting the model iteratively rewrite its own prompt, guided by feedback on prior outputs. ADO lin2025ado introduces DSP that imposes semantic constraints on iteratively proposed prompts to facilitate finding the optimal prompt. ProTeGi pryzant2023automatic generates natural language "corrections" that are applied as edits to the prompt, forming a textual analogue of gradient descent. PromptAgent wang2023promptagent casts prompt discovery as Monte-Carlo Tree Search, exploring instruction space strategically, while evolutionary approaches like PromptBreeder fernando2023promptbreeder maintain a population to discover increasingly effective instructions. REVOLVE zhang2024revolve further stabilizes long optimization runs by tracking the trajectory of model responses and applying smoothed updates. Pushing this autonomy to its limit, SPO xiang2025self creates a fully self-contained loop where the model generates its training data and uses pairwise preference comparison on its outputs to refine the prompt, eliminating the need for any external labeled data or human feedback. Collectively, these techniques demonstrate that an agent can autonomously improve its prompting policy, turning prompt text into a learnable component that co-evolves with the agent’s experience.
In complex systems, an agent often orchestrates a sequence of LLM calls or collaborates with other agents, making prompt design a multi-node problem. Frameworks such as DSPy represent an entire workflow as a graph whose sub-prompts are jointly tuned for a global objective khattab2023dspy . Trace wang2023trace , TextGrad yellamraju2024textgrad , and LLM-AutoDiff yin2025llm generalize this idea by treating each prompt as a parameter in a differentiable program and propagating natural-language “gradients” to refine every step. In collaborative scenarios, Multi-Agent System Search (MASS) zhou2025maasprompt first optimizes individual role prompts and then refines inter-agent communication patterns, while MAS-ZERO ke2025maszero dynamically proposes and revises role prompts to assemble an effective team for each new problem. Evolutionary systems such as EvoAgent yuan2024evoagent and AgentSquare shang2025agentsquare treat each agent along with prompts as the modules and use mutation and selection to discover specialized teams that outperform hand-crafted designs. These approaches extend PO from a single instruction to the language that defines whole workflows or societies of agents.
3.3 Tools
An agent’s capabilities are fundamentally defined by the tools it can wield. The trajectory of agent development is marked by a crucial evolution: from being mere tool users to becoming autonomous tool makers. This transition from relying on predefined, static toolsets to enabling agents to autonomously expand and refine their own skills is a critical leap towards cognitive self-sufficiency. This paradigm, where agents dynamically adapt their capabilities, allows them to solve a long tail of complex problems not envisioned by their initial designers. This evolution unfolds across three interconnected fronts: tool discovery, mastery, and management, as detailed in the subsections below.
Autonomous Discovery and Creation
The primary impetus for autonomous tool creation is to overcome the inherent limitations of a fixed toolset, granting agents the flexibility to innovate on demand. Methodologies for this now span a spectrum from opportunistic discovery to formalized synthesis. At one end, agents like Voyager build an ever-expanding library of skills through emergent trial-and-error, driven by an intrinsic motivation to explore complex, open-ended environments like Minecraft wang2023voyager . This exploratory approach is powerful for generating a wide array of skills but may lack precision. In contrast, systems like Alita and ATLASS take a more reactive approach, often employing retrieval-augmented generation (RAG) to search open-source code repositories or write new functions from scratch the moment a capability gap is identified qiu2025alita ; Haque2025ATLASS . At the other end of the spectrum lie highly structured frameworks that treat tool creation as a deliberate engineering process. CREATOR, for example, disentangles abstract tool creation (e.g., reasoning about the general structure of a reusable function for averaging temperatures over N days) from concrete tool usage (e.g., deciding how to apply that function to a specific city and time range), which enhances modularity and reusability qian2023creator . Even more formally, SkillWeaver analyzes successful human or agent task trajectories to propose, synthesize, and hone new skills into robust, reusable APIs, ensuring a higher degree of initial quality zheng2025skillweaver . Furthermore, frameworks like CRAFT demonstrate that creating specialized toolsets for specific domains is essential to complement general-purpose models, enabling expert-level performance without sacrificing adaptability yuan2024craft . RL-GPT liu2024rl integrates generated code implementations into the RL pipeline, leveraging these as tools to tackle complex tasks while addressing simpler ones directly using a Code-as-Policy approach. This integration dynamically adapts and evolves in response to environmental feedback, enabling continuous improvement. However, this burgeoning autonomy introduces significant challenges, particularly around safety and security. The unconstrained generation of code risks creating tools with exploitable vulnerabilities or unintended harmful behaviors, making automated verification and sandboxing critical areas for future research.
Mastery Through Iterative Refinement
The proliferation of self-created tools necessitates a robust mechanism for their mastery; a newly generated tool is often a brittle script, not a reliable function. This is where iterative refinement becomes essential. Frameworks like LearnAct and From Exploration to Mastery establish a critical self-correction loop where the agent learns from its own experience zhao2024learnact ; fromexplorationtomastery . This involves tackling the difficult "credit assignment" problem: determining precisely which line of code or which parameter was responsible for a failure. To do this, the agent analyzes a rich variety of feedback signals—including compiler errors, unexpected API return values, environmental state changes, or even implicit signals from a user’s subsequent actions. The goal is not only to debug the tool’s underlying code but also to refine its documentation (e.g., its docstring and argument descriptions), which is crucial for improving the agent’s ability to understand and correctly use the tool in the future. This refinement process also opens the door for valuable human-agent collaboration. While full autonomy is the ultimate goal, many systems can be designed with a "human in the loop," where a human expert can provide corrections, offer high-level suggestions, or validate a newly created tool. This collaborative approach can significantly accelerate the mastery process and ensure that the agent’s skills align with human intentions and safety standards. Ultimately, this self-honing process is what elevates a nascent skill into a dependable capability, ensuring the agent’s growing skill library increases not just in quantity, but more importantly, in quality and robustness.
Scalable Management and Selection
As an agent’s mastered skill library grows into the hundreds or thousands, it faces a "curse of abundance." The challenge shifts from creating tools to efficiently managing and selecting from them. A large library creates a massive search space, making traditional retrieval methods slow and inaccurate. To overcome this, ToolGen represents a fundamental paradigm shift by encoding tools as unique tokens within the language model’s vocabulary. This elegantly reframes tool retrieval as a generation problem, leveraging the transformer’s immense pattern-recognition capabilities to predict the most appropriate tool as a natural continuation of its thought process wang2025toolgen . Beyond selecting a single tool, advanced agents must also excel at tool composition—learning to chain multiple tools in novel sequences to solve multi-step problems. This is a higher-order management task. Architectural approaches like AgentSquare engage in a form of meta-learning, automatically searching the modular design space of an agent—including its planning, memory, and tool-use components—to find an optimal configuration for complex task execution shang2025agentsquare . As a logical endpoint to this evolutionary trend, visionary concepts like the Darwin Godel Machine propose a framework for open-ended evolution, where the agent can fundamentally rewrite its own core code. In this vision, the distinction between the agent and its tools blurs, leading to a recursive cascade of self-improvement that transcends tool enhancement alone zhang2025darwin . In essence, this entire evolutionary path aims to establish a closed and virtuous cycle: a truly autonomous agent that can perceive gaps in its capabilities, create novel solutions, master them through practice, and seamlessly integrate them into a coherently managed and ever-expanding repertoire.
3.4 Architecture
The defining feature of next-generation agentic systems is their intrinsic capacity for self-improvement. This marks a fundamental shift from systems with fixed capabilities to those that can autonomously enhance their performanceliu2025advances . By treating their own internal logic and collaborative structures as optimizable components, these systems can adapt their behavior and design in response to feedback, achieving a level of efficiency and effectiveness that static designs cannot match. This section details how this self-optimization is realized, first by examining improvements within single-agent systems and then by exploring the co-evolution of complex multi-agent systems.
3.4.1 Single-Agent System Optimization
LLM-Invoking Node Optimization
Optimizing a single LLM call is straightforward in isolation, but within an agentic system, it becomes a difficult credit assignment problem, as the effect of any single change is obscured by subsequent steps. Research addresses this by making node-level components optimizable, following two main strategies. The first focuses on refining nodes within a fixed agentic topology. A prime example is TextGrad yellamraju2024textgrad , which, inspired by backpropagation, uses "textual gradients" to propagate feedback from the final output backward through the workflow, guiding systematic, local refinements at each node without altering the system’s overall structure. The second, parallel strategy integrates this component-level optimization directly into the search for the system’s architecture itself. Under this approach, node characteristics become tunable parameters in a larger search space. For instance, frameworks can embed prompt engineering directly into the search loop, allowing the system to discover not just the optimal workflow but also the most effective instruction for each agent simultaneously zhou2025maasprompt . Similarly, EvoFlow zhang2025evoflow uses evolutionary algorithms to construct heterogeneous workflows by selecting the most suitable LLM for each task from a diverse pool. This holistic strategy enables the discovery of systems that are co-optimized for both their structure and individual agent capabilities, effectively balancing metrics like overall performance and cost ye2025xmas .
Autonomous-Agent Optimization
Building upon the optimization of individual LLM-invoking nodes, a more profound level of self-improvement targets the autonomous agent as a holistic entity. This evolution proceeds along two main fronts: optimizing the agent’s high-level architectural design and enabling the agent to directly modify its own source code. The first approach focuses on discovering the optimal agent structure. AgentSquare shang2025agentsquare exemplifies this by defining a modular design space of components like planners and memory modules, then using an evolutionary algorithm to find the most effective combination for a given task. The second front involves agents that dynamically rewrite their own operational code. This is seen in radical systems like the Darwin Gödel Machine zhang2025darwin , which recursively modifies its own Python codebase, and AlphaEvolve novikov2025alphaevolve , which uses evolutionary coding to improve specific algorithms. Similarly, Gödel Agent yin2025godelagentselfreferentialagent provides a self-referential framework for agents to analyze and alter their logic. Together, these two directions (optimizing the agent’s architectural “blueprint” and its functional code) demonstrate a key trend toward turning the agent’s fundamental structure and logic into learnable components.
3.4.2 Multi-Agent System Optimization
How agents are organized and communicate within a system (its topology) fundamentally determines its capacity for solving complex problems. The field has evolved from using fixed, human-designed communication structures to creating dynamic systems that automatically adapt their organization to a given task, allowing them to discover and exploit the most effective collaboration patterns. This evolution is explored along two major fronts: the optimization of static, explicit workflows and the co-evolution of dynamic, internal policies.
Agentic Workflow Optimization
The optimization of agentic workflows focuses on finding the most effective, often static, structure of communication and task delegation for a given problem. Early research established important foundations, with studies like AutoFlow li2024autoflow demonstrating the automated creation of linear workflows from natural language, and GPTSwarm zhuge2024gptswarm proposing a unifying graph-based framework. Concurrently, other foundational work explored how agents could evolve by using symbolic learning to distill their interaction experiences into an explicit, interpretable set of logical rules to guide future decisions zhou2024symbolic . This abstraction of systems into tunable components—whether nodes, edges, or symbolic rules—was crucial. However, these early systems often lacked a formal method for efficiently navigating the vast space of possible configurations and interactions.
The major breakthrough came when ADAS hu2024automated and AFlow zhang2024aflow formally defined this challenge as a search and optimization problem. ADAS set a theoretical vision by framing system design as a search through a Turing-complete space of code-based configurations. Building on this, AFlow made it practical by introducing reusable operators that represent common agentic patterns and by employing Monte Carlo Tree Search (MCTS) to efficiently navigate the enormous design space. Together, these works established a core methodology for treating agent system design as a tractable optimization problem, proving that automatically discovered workflows could outperform human-designed ones.
Following this formalization, research rapidly diversified toward creating customized agent systems for each specific query. Two primary strategies emerged: search-based and learning-based generation. Search-based methods, such as MaAS zhang2025maas , create a "supernet" of potential architectures and then sample a specialized system from it. In parallel, learning-based methods train models to generate effective topologies directly. ScoreFlow wang2025scoreflow , for instance, trains a generator using a novel preference optimization method, while FlowReasoner gao2025flowreasoner uses reinforcement learning to train a meta-agent that constructs a bespoke workflow on the fly. This line of query-specific generation continues to be an active area of research ye2025masgpt ; ke2025maszero . Furthermore, it is important to note that this process is not limited to the topology alone; many of these frameworks also perform node-level optimization in tandem, such as co-optimizing prompts or selecting heterogeneous models as an integral part of the architectural generation process zhang2024aflow ; zhou2025maasprompt ; zhang2025evoflow .
A key challenge for all search and learning methods is the computational cost of evaluating each potential workflow shang2025agentsquare . To address this, researchers have developed lightweight prediction models. Agentic Predictor trirat2025agentic is a prime example, training a model to accurately estimate a workflow’s performance based on its structural and semantic features without a full execution. By providing a fast and inexpensive evaluation proxy, these predictors significantly accelerate the optimization process, making the exploration of vast design spaces feasible zhang2025gnns .
Multi-Autonomous-Agent Optimization
Distinct from optimizing a system’s explicit workflow structure, this line of research focuses on how multiple autonomous agents can co-evolve their internal behavioral policies through interaction. This approach enables emergent capabilities like coordination, task delegation, and beneficial competition. For instance, ReMA wan2025remalearningmetathinkllms uses multi-agent reinforcement learning (MARL) to collaboratively train a high-level meta-thinker and a low-level executor, significantly improving performance on reasoning benchmarks. Building on this, GiGPO feng2025groupingrouppolicyoptimizationllm enhances MARL training by aggregating trajectories to provide more precise credit assignment, boosting success rates on long-horizon tasks. To support this direction, platforms like MARTI liao2025marftmultiagentreinforcementfinetuning provide open-source infrastructure for orchestrating and scaling the training of these language-model collectives. Collectively, these studies underscore multi-agent reinforcement learning as a promising route for cultivating group-level competencies unattainable by individual agents alone.
4 When to Evolve
The temporal dimension of self-evolution in LLM-based agents mainly concerns the relationship between learning processes and task execution. Therefore, the second key aspect of a self-evolving agent is identifying the evolving timing, i.e., at which stage the self-evolving strategy $f$ is invoked and applied to the agent system. To this end, we propose a taxonomy that distinguishes between two temporal modes of self-evolution: Intra-test-time self-evolution and inter-test-time self-evolution.
Intra-test-time self-evolution refers to adaptive processes that occur during task execution, where agents recognize their limitations on a specific problem and initiate targeted learning mechanisms to enhance their capabilities in real-time xi2024enhancing ; bi2024forest . This mode of evolution is characterized by its immediate coupling with the task at hand: the agent improves its problem-solving abilities for a specific problem encountered, creating a dynamic interplay between performance and adaptation.
Inter-test-time self-evolution refers to learning processes that occur between task completions, leveraging accumulated experiences to improve future performance. This category encompasses diverse methodological approaches: offline learning paradigms that extract knowledge from pre-collected datasets through iterative refinement zelikman2022star ; zelikman2024quiet , and online learning paradigms that continuously adapt based on streaming interaction data qi2024webrl ; qiu2025alita ; qian2024new ; wang2025mobile .
The implementation of self-evolution across these temporal phases leverages three fundamental learning paradigms in LLMs: in-context learning (ICL) dong2022survey ; min2021metaicl ; wies2023learnability , which adapts behavior through contextual examples without modifying parameters; supervised fine-tuning (SFT), which updates model weights through gradient-based optimization on labeled data devlin2018bert ; shen2024rethinking ; dong2023abilities ; and reinforcement learning (RL), which shapes behavior through reward-driven policy optimization kaelbling1996reinforcement ; sun2024llm ; zhang2025nemotron . While these learning paradigms remain conceptually consistent across temporal contexts, their instantiation differs in terms of data availability and learning objectives:
Intra-test-time is characterized by its online nature: learning data emerges dynamically during task execution, with optimization directly targeting performance enhancement on the immediate problem instance. This real-time coupling necessitates rapid adaptation mechanisms that can process learning data and feedback signals and modify behavior within the temporal constraints of active task-solving. On the other hand, inter-test-time is characterized by its retrospective nature: learning algorithms operate on historical data, whether from curated datasets or accumulated behavioral trajectories, with optimization objectives oriented toward improving expected performance across the task distribution rather than maximizing success on any specific problem instance. This temporal decoupling enables more sophisticated learning procedures that can identify cross-task patterns, consolidate diverse experiences, and develop generalizable capabilities without the immediacy constraints of active task execution.
Figure 5: An overview of when to evolve. The top pathway illustrates intra-test-time self-evolution, where adaptation (e.g., variant generation, verification, and policy update) occurs within task execution. The bottom pathway depicts inter-test-time self-evolution, where learning happens retrospectively through rollout, trajectory analysis, and policy updates.
4.1 Intra-Test-Time Self-Evolution
In intra-test-time self-evolution, agents engage in self-improvement processes that are intrinsically coupled with solving the immediate task at hand. The distinguishing characteristic of this temporal phase is its synchronous nature: feedback signals are generated and processed during task execution, with optimization objectives specifically targeted at improving performance on the current problem instance rather than generalizing to future tasks. Here, we introduce how the three learning paradigms are realized in this temporal phase.
In-Context Learning
Intra-test-time ICL methods leverage the model’s context window as a dynamic memory system for immediate adaptation without parameter modification. These approaches typically employ self-reflective mechanisms where agents analyze their own performance, generate verbal critiques or insights, and maintain these reflections in episodic memory buffers to guide subsequent decisions within the same task context shinn2023reflexion ; madaan2023self_refine . Some methods extend beyond simple reflection to include dynamic planning revision, where agents can modify their entire approach based on environmental feedback, switching between action execution and plan modification as needed. For instance, AdaPlanner sun2023adaplanner decomposes tasks into manageable sub-goals and predicts environmental feedback for each. During execution, its refiner component distinguishes between in-plan feedback (observations aligning with predictions) and out-of-plan feedback (deviating observations). For in-plan feedback, the refiner dynamically queries the LLM through a specialized ask_LLM() action to parse observations and extract pertinent information. For out-of-plan feedback, the refiner proactively revises the entire plan and resumes solving from an intermediate point, rather than restarting from scratch. This adaptive closed-loop framework eliminates the need for prior knowledge about feedback structures and enables more efficient decision-making. Similarly, TrustAgent hua2024trustagent employs rule-based plan revision during execution, modifying its approach based on language feedback to evolve toward safer planning strategies. These ICL methods demonstrate how test-time adaptation can achieve sophisticated behavioral modification without permanent model changes, maintaining flexibility while preserving the model’s general capabilities.
Supervised Fine-Tuning.
Intra-test-time SFT represents a paradigm shift where models perform immediate self-modification through learned meta-adaptation strategies. Self-adaptive language modeling zweiger2025self exemplifies this approach by generating “self-edits”, which are meta-level instructions that can restructure information representations, specify optimization hyperparameters, or invoke tools for data augmentation and gradient computation. These self-edits trigger immediate supervised fine-tuning, resulting in persistent weight updates that adapt the model to the current task. The key innovation lies in the meta-learning phase, where reinforcement learning trains models to produce effective self-edits by using the downstream performance of the updated model as the reward signal, essentially teaching models how to teach themselves.
Reinforcement Learning.
Intra-test-time RL enables models to develop new capabilities on-demand when encountering problems beyond their current competence. LADDER simonds2025ladder demonstrates this through its test-time reinforcement learning (TTRL) mechanism: upon identifying a particularly challenging problem, the system generates a focused set of related problem variants and conducts intensive, targeted reinforcement learning specifically for that problem class. This approach transforms insurmountable challenges into learning opportunities, allowing models to expand their problem-solving repertoire during deployment rather than failing or providing suboptimal solutions. The method represents a form of just-in-time skill acquisition, where computational resources are invested precisely when and where they are needed most.
4.2 Inter-Test-Time Self-Evolution
Inter-test-time self-evolution represents the predominant learning process in autonomous agents, wherein adaptation occurs following task execution rather than during it. In this temporal mode, agents complete a given task, extract feedback signals, including explicit rewards gao2024designing , gradients amari1993backpropagation ; bottou2010large , and performance metrics ge2023openagi , and subsequently leverage this information to enhance their capabilities for future problem-solving. This retrospective learning process decouples task performance from capability improvement, allowing agents to consolidate experiences, identify patterns of success and failure, and systematically refine their behavioral policies without the computational constraints imposed by real-time task demands.
In-Context Learning.
Inter-test-time in-context learning has emerged as a widely adopted approach for agent self-improvement. This paradigm leverages execution results and feedback from previous tasks as contextual information for future problem-solving. Wang et al. wang2024agent demonstrate this principle by inducing workflows from agent action histories and incorporating them into the context for subsequent tasks. The field of in-context reinforcement learning (ICRL) moeini2025survey ; laskin2022context ; lee2023supervised extends this concept by maintaining histories of observations and actions within the agent’s context window. These methods exploit the hypothesis that pre-trained neural networks can implement implicit reinforcement learning algorithms within their forward pass, processing contextual information to adapt behavior without parameter updates kirsch2023towards . A defining characteristic of ICRL is in-context improvement: the phenomenon whereby agent performance progressively enhances as task-relevant information accumulates in the context, enabling sophisticated adaptation through attention mechanisms rather than gradient-based learning.
Supervised Fine-Tuning.
Inter-test-time SFT chen2025sft methods establish a paradigm of iterative self-improvement through synthetic data generation and self-evaluation. SELF lu2023self pioneered meta-cognitive training, where models first acquire self-feedback and self-refinement capabilities, then iteratively generate responses to unlabeled instructions and enhance them through self-critique. STaR zelikman2022star and Quiet-STaR zelikman2024quiet focus on reasoning improvement through rationalization—models attempt problems, then generate explanations for correct answers they initially failed to solve, creating augmented training data that combines successful attempts with post-hoc reasoning. SiriuS zhao2025sirius extends this to sequential problem-solving, maintaining repositories of correct solutions while augmenting failures through multi-stage refinement involving feedback incorporation, regeneration, and rephrasing. These methods share a core insight: models can bootstrap their own improvement by learning to evaluate and enhance their outputs, creating high-quality training signals from initially imperfect attempts without extensive human supervision.
Reinforcement Learning.
Inter-test-time RL leverages unconstrained computational resources to optimize agents through extensive environmental interaction and sophisticated curriculum design. RAGEN wang2025ragen and DYSTIL wang2025dystildynamicstrategyinduction employ online reinforcement learning for multi-turn interactive tasks, continuously refining policies through on-policy learning in simulated dialogues. Learning Like Humans zhang2025learning introduces cognitive-inspired training with adaptive difficulty progression, combining on-policy exploration with off-policy efficiency and expert demonstrations to accelerate learning. Domain-specific applications demonstrate the versatility of inter-test-time RL: WebRL qi2024webrl develops web navigation agents through self-evolving curricula that automatically adjust task complexity based on performance, while DigiRL bai2024digirl enables device-control agents to master in-the-wild interactions through autonomous reinforcement learning. These approaches exploit the pre-deployment phase to engage in extensive trial-and-error learning, developing robust policies through thousands of interactions that would be impractical during real-time deployment.
5 How to Evolve
Table 3: Overview of Reward-based, Imitation/Demonstration, and Population-based Learning Methods for Self-Evolving Agents. This table categorizes key approaches based on the following criteria: (1) Feedback Type: the type of feedback used, including language-based rationales and numerical rewards. (2) Feedback Source: the origin of the feedback, either internal (model-generated) or external (provided externally). (3) Learning Method: the learning paradigm applied, such as in-context learning (ICL), supervised fine-tuning (SFT), reinforcement learning (RL), and evolutionary algorithms; (4) Updated Components: which parts of the model are updated, either full parameters or a subset of the model. (5) Update Timing: the stage during the agent’s evolution when updates are applied, such as pre-training, pre-test, or test-time.
Method Feedback Type Feedback Source Learning Method Updated Components Update Timing
Reward-based Evolution Methods
Reflexion shinn2023reflexion language internal ICL context test-time
AdaPlanner sun2023adaplanner language external + internal ICL context test-time
AgentS2 agashe2025agents2 language external ICL context test-time
SELF lu2023self language external + internal SFT full params pre-test time + test-time
SELF-REFINE madaan2023self_refine language internal ICL context test-time
SCoRe kumar2024training numerical external RL full params pre-test time
PAG jiang2025pag numerical external RL full params pre-test time
TextGrad yellamraju2024textgrad language external ICL context pre-test time / test-time
SRSI simonds2025selfrewardingselfimproving language internal RL full params pre-test time
Self-Train LM shafayat2025largereasoningmodelsselftrain numerical internal RL full params pre-test time
MM-UPT wei2025unsupervisedposttrainingmultimodalllm numerical internal RL full params pre-test time
CoVo zhang2025consistentpathsleadtruth numerical internal RL full params pre-test time
SWE-agent du2025swedevevaluatingtrainingautonomous language external ICL context test-time
SICA robeyns2025self numerical external ICL codebase(tools, workflows, prompts) test-time
Feedback Friction jiang2025feedbackfriction language external ICL context test-time
USEagent applis2025useagent language external ICL context test-time
DYSTIL wang2025dystildynamicstrategyinduction language + numerical external + internal SFT+RL full params pre-test time + test-time
OTC-PO wang2025otcpo numerical external RL full params pre-test time
AUTORULE wang2025autorule language + numerical external + internal RL full params pre-test time
EGSR zhang2025learning numerical external RL full params pre-test time
LADDER simonds2025ladder numerical external RL full params pre-test time
RAGEN wang2025ragen numerical external RL full params test-time
SPIRAL liu2025spiral numerical internal RL full params pre-test time
ICRL Prompting song2025reward numerical external + internal RL full params test-time
MATH-SHEPHERD wang2023math numerical external RL full params pre-test time
AgentPRM choudhury2025process numerical external SFT+RL full params pre-test time
Agent Q putta2024agentq numerical external RL full params pre-test time
GiGPO feng2025group numerical external RL full params pre-test time
SPA-RL wang2025spa numerical external RL full params pre-test time
Self-Instruct wang2022self language internal SFT full params pre-test time
WizardLM xu2024wizardlm language internal SFT full params pre-test time
OS-Genesis sun2024genesis numerical external SFT full params pre-test time
UI-Genie xiao2025ui numerical external SFT partial params pre-test time
GUI-R1 luo2025gui numerical external SFT+RL full params pre-test time
InfiGUI-R1 liu2025infigui numerical external SFT+RL full params pre-test time
Voyager wang2023voyager language external ICL context test-time
SwiftSage lin2023swiftsage language external ICL context test-time
AutoWebGLM lai2024autowebglm language external SFT+RL full params pre-test time
DigiRL bai2024digirl language external RL partial params pre-test time
WebRL qi2024webrl language external SFT+RL full params pre-test time
Let’s Verify Step-by-Step lightman2023let language external SFT full params pre-test time
AlphaMath chen2024alphamath numerical external SFT full params pre-test time
rStar-Math guan2025rstar numerical external SFT full params pre-test time
DistRL wang2024distrl language external RL full params pre-test time + test-time
MobileGUI-RL shi2025mobileguirladvancingmobilegui language external RL full params pre-test time
Imitation and Demonstration Learning Methods
STaR zelikman2022star language + numerical internal SFT full params pre-test time
V-STaR hosseini2024v numerical external + internal SFT + RL partial params pre-test time
AdaSTaR koh2025adastar numerical internal SFT full params pre-test time
STIC deng2024enhancing language internal RL + SFT partial params pre-test time
GENIXER zhao2024genixer language external SFT full params pre-training
SiriuS zhao2025sirius language + numerical internal SFT full params pre-test time
SOFT tang2025bridging language internal SFT not specified pre-test time
RISE qu2024recursive language + numerical internal + external SFT full params pre-test time
IoE li2024confidence numerical internal / / test-time
Population-based and Evolutionary Methods
DGM zhang2025darwin numerical external ICL codebase (tools, workflows, prompts) test-time
EvoMAC hu2024self language external ICL team composition, workflow, prompts test-time
SPIN chen2024self language internal RL full params pre-test time
GENOME zhang2025nature numerical external Evolution Alg. partial params pre-test time
SPC chen2025spc numerical internal SFT+RL critic params pre-test time + test-time
Puppeteer dang2025multi numerical external RL planner policy pre-test time / between tasks
MedAgentSim almansoori2025self language external ICL context (knowledge base) test-time
STL mendes2025language language + numerical internal SFT value model pre-test time
MDTeamGPT chen2025self language external ICL context (knowledge base) test-time
The pursuit of self-evolution lies at the heart of building advanced, autonomous, and increasingly general artificial intelligence. For large language models (LLMs) and their agentic extensions, the question of how to continually, autonomously, and efficiently evolve their capabilities has become a central challenge. Therefore, the third key aspect of a self-evolving agent is to instantiate an effective evolving strategy $f$, i.e., how to transform an agent system $\Pi=(\Gamma,\{\psi_{i}\},\{C_{i}\},\{\mathcal{W}_{i}\})$ to its new state $\Pi^{\prime}=(\Gamma^{\prime},\{\psi^{\prime}_{i}\},\{C^{\prime}_{i}\},\{\mathcal{W}^{\prime}_{i}\})$. Unlike traditional approaches that rely on static datasets or one-time supervised fine-tuning, self-evolution emphasizes an ongoing process where models learn from real-world interactions, actively seek feedback, self-reflect, generate or curate new data, and adapt their strategies in response to dynamic environments. This continuous evolution is not merely a matter of scaling up data or computation; it requires the agent to acquire a spectrum of meta-capabilities, including self-correction, autonomous data generation, knowledge transfer, and multi-agent collaboration. As a result, the landscape of self-evolution has become increasingly rich and multi-faceted, with each methodological branch exploring different axes of feedback, learning paradigms, data sources, and evolutionary scales.
This chapter aims to systematically map and analyze the major families of self-evolution methods, providing a unified framework for understanding their principles, mechanisms, and interactions. We begin with reward-based evolution, which centers on the design of reward signals—ranging from natural language feedback and internal confidence metrics to external or implicit signals—to guide iterative self-improvement. Next, we examine imitation and demonstration learning, where agents improve by learning from high-quality exemplars, either self-generated or provided by other agents or external sources. This paradigm is particularly powerful when demonstrations are abundant or can be autonomously synthesized, and it has driven significant progress in both reasoning and multimodal domains. Finally, we introduce population-based and evolutionary methods, which draw inspiration from biological evolution and collective intelligence. These approaches maintain populations of agent variants or collaborating agents, leveraging mechanisms such as selection, mutation, crossover, and competition to explore the solution space in parallel, foster diversity, and enable the emergence of novel strategies or architectural innovations.
5.1 Reward-based Self-Evolution
The capacity for self-improvement is a cornerstone of advanced intelligence. In the context of Large Language Models (LLMs), this manifests as a dynamic process of reward-driven evolution, where models iteratively learn from their own outputs and interactions to refine their capabilities. The design of the reward signal, which serves as the guiding feedback, is crucial; it determines the nature, efficiency, and effectiveness of the learning process. In this section, we systematically review the main methodologies for reward design, categorized by the nature of the feedback: textual feedback, internal confidence, external rewards, and implicit rewards.
Figure 6: Overview of reward-based self-evolution strategies, categorized into textual, implicit, internal, and external rewards, each associated with distinct feedback sources and mechanisms.
Textual Feedback
Textual Feedback leverages the native modality of LLMs—natural language—to provide detailed, interpretable instructions for refinement. Unlike scalar rewards, textual feedback encapsulates nuanced critiques and actionable suggestions. Recent frameworks such as Reflexion (shinn2023reflexion, ), AdaPlanner (sun2023adaplanner, ), AgentS2 (agashe2025agents2, ), SELF (lu2023self, ), Self-Refine (madaan2023self_refine, ), SCoRe (kumar2024training, ), PAG (jiang2025pag, ), and TextGrad (yellamraju2024textgrad, ) exemplify this direction. For instance, Reflexion proposes “verbal reinforcement learning,” where agents reflect in natural language on their past trials, storing these reflections as episodic memory to guide future decisions. AdaPlanner enables closed-loop adaptive planning by allowing LLM agents to revise their plans based on both in-plan and out-of-plan feedback, while also mitigating hallucination via code-style prompts and leveraging skill discovery. Self-Refine and SELF further explore iterative self-feedback and self-correction, demonstrating that even state-of-the-art models can be improved via multi-turn, language-based self-critique, without additional supervised data or external reinforcement. Such frameworks highlight the power of language as a reward channel, enabling nuanced, flexible, and sample-efficient self-improvement.
Internal Rewards
Internal Confidence-based rewards move away from external signals and instead exploit internal metrics such as the model’s probability estimates or certainty. This paradigm leverages the model’s intrinsic understanding to guide improvement without relying on external supervision. Methods such as Confidence-Informed Self-Consistency (CISC) (taubenfeld2025cisc, ), Self-Ensemble (xu2025selfensemblemitigatingconfidencedistortion, ), Self-Rewarding Self-Improving (simonds2025selfrewardingselfimproving, ), scalable best-of-N selection via self-certainty (kang2025scalablebestofnselectionlarge, ), and Self-Rewarding Language Models (yuan2025selfrewardinglanguagemodels, ) allow models to self-evaluate and calibrate their responses based on internal confidence metrics. For example, CISC weights reasoning paths by confidence scores to improve both accuracy and computational efficiency, effectively filtering high-quality solutions from multiple candidates. Self-Ensemble mitigates confidence distortion by dividing choices into smaller, more manageable groups and aggregating predictions to reduce overconfidence bias. Self-Rewarding Language Models demonstrate that models can act as their own reward function, generating training data through self-instruction and self-evaluation cycles. These approaches can reduce reliance on human labels and external evaluators, enabling scalable and autonomous self-improvement loops that can operate continuously without human intervention.
External Rewards
External Rewards are derived from sources outside the model, such as the environment, majority voting, or explicit rules. Majority voting (shafayat2025largereasoningmodelsselftrain, ; wei2025unsupervisedposttrainingmultimodalllm, ; zhang2025consistentpathsleadtruth, ) uses consensus among multiple model outputs as a proxy for correctness, providing a self-generated but grounded reward signal. Environment feedback, including tool-based signals, is central to agentic LLM research (e.g., SWE-Dev (du2025swedevevaluatingtrainingautonomous, ), SICA (robeyns2025self, ), Feedback Friction (jiang2025feedbackfriction, ), USEagent (applis2025useagent, ), DYSTIL (wang2025dystildynamicstrategyinduction, )), where agents learn through direct interaction with real-world environments and tools. Rule-based rewards (wang2025otcpo, ; wang2025autorule, ; zhang2025learning, ; simonds2025ladder, ; wang2025ragen, ; liu2025spiral, ) use explicit constraints or logical rules as verifiable signals, particularly effective in the domains of mathematical reasoning, game play, and structured problem solving. These methods offer objective, reliable supervision but may require significant engineering or be limited in expressiveness.
Implicit Rewards
Implicit Reward frameworks hypothesize that LLMs can learn from feedback signals even when not explicitly labeled as rewards. For instance, “Reward Is Enough” (song2025reward, ) demonstrates that LLMs can perform in-context reinforcement learning using simple scalar signals embedded in the context window, improving their responses over rounds without explicit RL fine-tuning or supervision. This reveals an inherent capacity for models to interpret and learn from implicit feedback cues present in their input context. Recent work has expanded this concept by showing that LLMs inherently encode reward-like signals through their standard training objectives. Endogenous reward (li2025generalist, ) reveal that standard next-token prediction implicitly learns a generalist reward function, which can be extracted from model logits without additional training. Moreover, ImPlicit Self-ImprovemenT (PIT) framework (wang2024enabling, ) implicitly learns the improvement goal from human preference data without extra human efforts by maximizing the quality gap of the response conditioned on a reference response. Unlike rule-based or environment-derived external rewards, implicit reward methods offer unique advantages by discovering and utilizing reward signals that are inherently present in language modeling.
5.2 Imitation and Demonstration Learning
Imitation and demonstration learning is a paradigm in which self-evolving agents improve their capabilities by learning from high-quality exemplars, which may be generated by the agents themselves, other agents, or external sources. Unlike reward-based methods that rely on explicit reward signals, imitation-based approaches focus on reproducing and refining successful behavioral patterns through iterative self-training and bootstrapping mechanisms. This approach is particularly effective when high-quality demonstrations are available or can be autonomously generated, allowing agents to bootstrap their capabilities with minimal external supervision.
5.2.1 Self-Generated Demonstration Learning
Self-generated demonstration learning involves agents creating their own training data through iterative refinement processes, where the models learn to improve by generating and selecting high-quality examples from their own outputs.
Bootstrapping Reasoning Capabilities. zelikman2022star introduces the foundational framework for self-generated demonstration learning, enabling language models to bootstrap their reasoning capabilities through iterative self-training. This process involves generating reasoning chains for problems, fine-tuning on correct solutions, and repeating this cycle to progressively improve performance without the need for ground-truth reasoning paths. Building on this framework, recent advancements have refined the bootstrapping process through more sophisticated training strategies. For instance, hosseini2024v proposes a verifier-guided self-training approach, where separate verifier models assess the quality of generated reasoning chains before they are incorporated into the training data, enhancing the reliability of self-improvement. Additionally, koh2025adastar introduces adaptive data sampling strategies that dynamically adjust the composition of training data based on model performance across various reasoning tasks, thereby mitigating overfitting to specific problem types.
Multimodal Self-Training. Extending self-training to multimodal domains presents unique challenges in generating high-quality demonstrations that span both visual and textual modalities. deng2024enhancing demonstrates how vision-language models can improve iteratively by training on their own generated image descriptions and visual reasoning chains. The approach leverages the model’s existing visual understanding to generate detailed image descriptions, which are subsequently used to fine-tune the model’s visual perception in a bootstrapping manner. zhao2024genixer builds on this concept by empowering multimodal large language models to serve as powerful data generators, producing diverse training examples across different modalities and tasks through advanced prompt engineering and quality filtering mechanisms.
5.2.2 Cross-Agent Demonstration Learning
Cross-agent demonstration learning involves agents learning from demonstrations provided by other agents, either within the same system or from external sources, enabling knowledge transfer and collaborative improvement.
Multi-Agent Bootstrapped Reasoning. zhao2025sirius presents a framework for multi-agent systems to learn from each other’s successful demonstrations through bootstrapped reasoning. The system maintains an experience library containing successful interaction trajectories generated by different agents, facilitating efficient knowledge sharing and collaborative improvement. Each agent can leverage the collective experience of the entire system, thereby accelerating the learning process and enabling the discovery of diverse solution strategies. This framework illustrates how agents can specialize in different aspects of complex tasks while benefiting from the accumulated knowledge of the entire system.
Domain-Specific Demonstration Learning. Domain-specific applications of demonstration learning have proven especially effective in specialized fields where expert knowledge can be effectively transferred through demonstrations. In recommendation systems, techniques such as self-optimized fine-tuning tang2025bridging enable LLM-based recommender systems to learn from their own successful recommendation patterns, creating a feedback loop that enhances personalization over time. The system generates high-quality recommendation demonstrations from successful user interactions and uses these to fine-tune the underlying language model, ultimately leading to more accurate and personalized recommendations.
5.2.3 Hybrid Demonstration Learning
Hybrid demonstration learning combines both self-generated and external demonstrations to create more robust and diverse training regimens that leverage the strengths of each approach.
Recursive Self-Improvement. qu2024recursive demonstrates how agents can be trained to systematically improve their behavior through structured self-reflection and demonstration generation. This approach enables language model agents to introspect on their reasoning processes, identify areas for improvement, and generate corrective demonstrations to address these weaknesses. This recursive process establishes a continuous improvement loop, where agents become increasingly skilled at self-diagnosis and self-correction, leading to more robust and adaptable behavior.
Confidence-Guided Demonstration Selection. Recent developments have focused on more sophisticated mechanisms for selecting high-quality demonstrations from both self-generated and external sources. Confidence-based approaches li2024confidence utilize the model’s uncertainty estimates to determine which demonstrations are most likely to contribute positively to learning, filtering out potentially detrimental or low-quality examples. This method addresses a critical challenge in demonstration learning: poor-quality demonstrations can degrade performance. By ensuring that only high-confidence, high-quality examples are used for training, this approach helps to maintain the integrity of the learning process.
The effectiveness of imitation and demonstration learning approaches is highly dependent on the quality and diversity of the available demonstrations. While these methods can yield impressive results when high-quality exemplars are present, they face challenges in domains where good demonstrations are scarce or where the optimal behavior is not well-represented in the available data. Future research directions include developing more sophisticated demonstration selection and generation strategies, improving the robustness of learning from imperfect demonstrations, and creating better mechanisms for combining demonstrations from multiple sources.
5.3 Population-based and Evolutionary Methods
Population-based and evolutionary methods represent a fundamentally different paradigm for agent evolution compared to the reward-based and imitation-based approaches discussed in previous sections. While reward-based methods typically optimize individual agents through iterative reward signals and imitation learning relies on learning from demonstrations, population-based methods draw inspiration from biological evolution and collective intelligence. These approaches maintain multiple agent variants simultaneously, allowing for parallel exploration of the solution space and the emergence of diverse capabilities through mechanisms such as selection, mutation, crossover, and competitive interaction zhang2025nature . This enables broader search coverage and the discovery of novel solutions that might be missed by gradient-based optimization. This approach is particularly valuable when the solution space is complex, multimodal, or when the optimal strategy requires fundamental architectural changes rather than parameter fine-tuning.
5.3.1 Single Agent Evolution
Single-agent evolutionary approaches focus on evolving individual agents through population-based mechanisms, where multiple variants of an agent compete and evolve over time. These methods can be broadly categorized into two main paradigms: learning from evolution and self-play from multiple rollouts.
Learning from Evolution. This paradigm draws directly from biological evolution, maintaining populations of agent variants and applying evolutionary operators to discover improved capabilities. The Darwin Gödel Machine (DGM) zhang2025darwin exemplifies this approach through open-ended evolution of self-improving agents that maintain an archive of all historical versions, enabling branching from any past "species" rather than linear optimization. The system achieves self-referential improvement by allowing agents to directly modify their own Python codebase, with evolution driven by empirical performance on coding benchmarks and parent selection balancing performance scores with novelty rewards for diverse exploration. Complementing this code-level evolution, the Nature-Inspired Population-Based Evolution (GENOME) framework zhang2025nature directly applies genetic algorithms to language model parameter evolution, maintaining populations and using crossover, mutation, and selection operators on model weights. GENOME+ extends this with particle swarm optimization concepts, adding inheritance mechanisms and ensemble methods that demonstrate gradient-free evolutionary optimization can effectively improve model capabilities through parameter space exploration.
Self-Play from Multiple Rollouts. This paradigm focuses on agents improving through iterative self-competition and rollout-based learning, where agents generate multiple trajectories and learn from their own exploration. Self-Play Fine-Tuning (SPIN) chen2024self establishes the foundation by having current models compete against previous versions, creating evolutionary pressure where only improving strategies survive without external annotations. SPC chen2025spc advances this through sophisticated adversarial co-evolution, where a "sneaky generator" learns to create deceptive errors while a "step critic" evolves to detect increasingly subtle mistakes, using automated validation to sustain improvement without human step-level annotations. STL mendes2025language demonstrates self-teaching evolution through iterative lookahead search, where value models generate training data from their own exploratory rollouts, combining numerical value learning with natural language reasoning chains to bootstrap continuous improvement. These approaches share the principle of using agents’ own generated experiences as learning signals, creating self-sustaining improvement cycles that evolve without external supervision.
5.3.2 Multi-Agent Evolution
Multi-agent evolutionary methods extend population-based approaches to evolving entire teams or networks of agents, focusing on optimizing collective behavior, coordination strategies, and collaborative architectures. These approaches can be categorized into two main paradigms based on their evolution mechanisms: System Architecture Evolution and Knowledge-Based Evolution.
System Architecture Evolution. This paradigm focuses on evolving the structural and coordination aspects of multi-agent systems, including team composition, orchestration strategies, and workflow optimization. EvoMAC hu2024self introduces a framework that mimics neural network training for multi-agent systems, implementing "textual backpropagation" where compilation errors and test failures serve as loss signals to drive iterative modifications of agent team composition and individual prompts. A specialized "updating team" analyzes textual feedback to identify problematic agents and generate modification instructions, effectively implementing gradient-based optimization in the space of agent configurations rather than model parameters. Building on this structural evolution concept, Puppeteer dang2025multi takes a different approach by focusing on coordination strategy evolution rather than team composition changes. The system employs a centralized orchestrator that evolves its decision policy through reinforcement learning, dynamically selecting which agents to activate at each step while balancing task performance with computational cost. This "puppeteer-puppet" paradigm demonstrates how architectural evolution can occur at the coordination level, discovering efficient collaboration patterns and emergent behaviors such as tighter coordination among core agents and sophisticated cyclic interaction patterns.
Knowledge-Based Evolution. This paradigm emphasizes evolving the collective knowledge and experience of multi-agent teams through memory accumulation and case-based learning, primarily operating through in-context learning or in-context-like adaptation rather than parameter updates. MDTeamGPT chen2025self establishes the foundation for this approach through a dual knowledge base system, implementing CorrectKB for storing successful cases and ChainKB for capturing failure reflections, enabling the system to learn from both successes and mistakes through structured case retrieval and reasoning enhancement. Extending this medical consultation framework, MedAgentSim almansoori2025self demonstrates how such knowledge-based evolution can be applied to real-world diagnostic scenarios, accumulating experience from patient interactions and using retrieval-augmented generation to improve consultation quality over time. PiFlow pu2025piflow applies this paradigm to scientific discovery, maintaining a trajectory of principle-outcome pairs and using them to steer hypothesis generation through information-theoretical optimization.
5.4 Cross-cutting Evolutionary Dimensions
Figure 7: Illustration of cross-cutting evolutionary dimensions in agent self-evolution, structured along three key axes: learning paradigm (offline/online), policy consistency (on/off-policy), and reward granularity (process-based, outcome-based, and hybrid). These dimensions jointly characterize how autonomous agents generate data, interact with environments, adapt policies, and receive feedback, providing a structured lens for analyzing reward-based, imitation-based, and population-based evolution strategies.
Agent self-evolution is a multifaceted process characterized by a number of cross-cutting dimensions that shape how agents learn, adapt, and improve over time. Beyond any single learning algorithm or supervision signal, these dimensions define the core principles underlying the design and analysis of autonomous agents. In this section, we systematically compare the major families of self-evolution methods—reward-based, imitation/demonstration-based, and population-based—along several key axes, such as learning paradigm (online vs. offline), policy consistency (on-policy vs. off-policy), and reward granularity (process-based, outcome-based, or hybrid). We further highlight additional dimensions, including feedback types, data sources, sample efficiency, stability, and scalability, as summarized in Table 4. This comprehensive comparison provides a unified perspective for understanding the strengths, limitations, and design trade-offs inherent in different approaches to agent evolution.
Table 4: Comparison of self-evolution method families along key dimensions.
Dimension Reward-based Imitation/Demonstration Population-based
Feedback Type Scalar reward, natural language, confidence, external signals Demonstration trajectories, exemplars, rationales Fitness scores, task success, competitive signals
Data Source Self-generated, environment, external rules Self-generated or other agents, humans Population generations, multi-agent systems
Reward Granularity Outcome/process/hybrid (flexible) Usually outcome/process (via demo steps) Often outcome-level, sometimes process via competition
Online/Offline Both (reward learning, RL, DPO, SFT) Typically offline, sometimes online demo mining Online evolution or batch population updates
On/Off-policy Both (DPO, Reflexion, GRPO) Primarily off-policy, but online variants can be on-policy Off-policy (population); self-play is on-policy
Sample Efficiency Moderate (depends on reward sparsity) High (if demo quality is high) Usually low (needs many trials)
Stability Sensitive to reward design Sensitive to demo quality/diversity Sensitive to population size/diversity
Scalability Good with automation Limited by demo collection High but resource-intensive
5.4.1 Online and Offline Learning
Another fundamental dimension in the design of self-evolving agents is the learning paradigm, which can be broadly categorized as either offline or online. This distinction depends on whether the agent’s evolutionary updates are performed on a static, pre-collected dataset of experiences (offline) or through continuous, direct interaction with a live environment (online).
Offline Learning In the offline learning paradigm, the learning phase is decoupled from live task execution. The offline process typically involves cycles of offline data generation, filtering, and model fine-tuning, focusing on building a powerful and generalist foundational model before deployment. A primary strategy in this domain is LLM bootstrapping, where a model enhances its own capabilities using its self-generated content. For example, Self-Instruct wang2022self shows how a language model can bootstrap its own instruction-following ability by generating new instructions, paired with its own responses, creating a synthetic dataset for fine-tuning. Building on this, WizardLM xu2024wizardlm demonstrates how to progressively evolve the complexity of these self-generated instructions, pushing the model’s capabilities on more challenging tasks. In the context of GUI and Web agents, offline learning often involves leveraging pre-collected high-quality trajectories for supervised fine-tuning (SFT). OS-Genesis sun2024genesis introduced a reverse task synthesis method for automatic trajectory creation. Similarly, UI-Genie xiao2025ui employs a unified reward model for trajectory evaluation and a self-improving loop to generate high-quality trajectories iteratively. Both approaches focus on curating a rich SFT dataset to enhance the agent’s capabilities to solve complex tasks. Beyond SFT, offline methods also incorporate reinforcement learning performed on a static dataset of agent-environment interactions. For example, GUI-R1 luo2025gui and InfiGUI-R1 liu2025infigui utilize rule-based rewards and apply R1-style guo2025deepseek training on offline GUI datasets.
Online Learning In contrast, online learning enables an agent to learn and adapt continuously while it interacts with a live or simulated environment. Feedback from each action is used to update the agent’s policy, plan, or knowledge base in real-time. This allows for greater adaptability to dynamic or unseen situations. Some agents evolve online not by updating their model weights, but by refining their plans and skill libraries on the fly. For example, Voyager wang2023voyager presents an LLM-powered agent that learns to play Minecraft by continuously exploring, generating its own curriculum of tasks, and building a persistent skill library from direct experience. AdaPlanner sun2023adaplanner focuses on adapting its plan within a task; it generates an initial plan, receives feedback from the environment, and refines the plan online. Similarly, SwiftSage lin2023swiftsage operates with a fast-and-slow thinking process, where it can reflect on failures of its fast, intuitive mode and switch to a more deliberate, tool-using slow mode, adapting its strategy online based on task difficulty. Reinforcement Learning serves as a fundamental mechanism for online learning, enabling agents to learn from environmental reward signals. DigiRL bai2024digirl demonstrates how to train device-control agents in the wild using autonomous RL, while DistRL wang2024distrl proposes an asynchronous distributed framework to make such on-device training feasible. MobileGUI-RL shi2025mobileguirladvancingmobilegui addresses the specific challenges of training GUI agents in online mobile environments by introducing a synthetic task generation pipeline combined with group relative policy optimization (GRPO) through trajectory-aware rewards.
5.4.2 On-policy and Off-policy Learning
While the previous section examined the timing of data collection and learning (online vs offline), this section focuses on the policy consistency aspect of agent evolution - specifically, whether agents learn from experiences generated by the same policy they are trying to improve (on-policy) or from experiences generated by different policies (off-policy). This distinction is crucial for understanding how agents utilize their experiential data and manage the trade-offs between learning stability and sample efficiency during the evolutionary process.
On-policy Learning. On-policy approaches require agents to learn exclusively from experiences generated by their current policy, ensuring policy consistency but often at the cost of sample efficiency. Reflexion shinn2023reflexion exemplifies this approach through its iterative self-reflection mechanism. The agent generates responses using its current policy, receives feedback on failures, and immediately incorporates this feedback to update its reasoning process for the next iteration. GRPO shao2024deepseekmath and DAPO yu2025dapo continue this path and show the effectiveness of multiple rollouts. The agent always learns from its current behavior, maintaining strict policy consistency. In agent settings, on-policy methods provide excellent learning stability and avoid distribution mismatch issues that plague off-policy methods. However, they suffer from low sample efficiency, as each policy update requires fresh data collection, making them computationally expensive for complex multi-step reasoning or tool use scenarios where generating high-quality trajectories is costly.
Off-policy Learning. Off-policy approaches allow agents to learn from experiences generated by different policies, including previous versions, other agents, or human demonstrations, significantly improving sample efficiency at the cost of potential distribution mismatch. yuan2401self demonstrates a sophisticated off-policy approach where model $M_{t+1}$ learns from preference data generated by the previous version $M_{t}$. The system handles distribution shift through DPO’s built-in KL divergence constraint with the reference policy, preventing the new policy from deviating too far from the data-generating policy. yuan2023rrhf showcases another powerful off-policy paradigm by learning from diverse response sources—including other models, humans, and different sampling strategies—through ranking-based supervision. The method elegantly sidesteps distribution shift by treating alignment as a ranking problem rather than requiring policy consistency. zhao2025sirius illustrates off-policy learning in multi-agent settings, where agents learn from an "experience library" containing successful interaction trajectories generated by previous policy versions, enabling efficient reuse of expensive multi-agent coordination data. In agent settings, off-policy methods excel in sample efficiency, allowing agents to leverage historical data, expert demonstrations, and cross-agent learning. They are particularly valuable for multi-step reasoning where successful trajectories are rare and expensive to generate, and for tool use scenarios where agents can learn from diverse execution examples without repeated environmental interaction. However, they face challenges with distribution shift, reward hacking (where agents exploit inconsistencies between training and deployment policies), and the need for careful regularization to maintain training stability.
5.4.3 Reward Granularity
Another critical choice in the reward design is its granularity, which determines at what level of detail the agent receives its learning signal. Reward granularity ranges from coarse-grained outcome-based rewards, which evaluate the overall task completion, to fine-grained process-based rewards that assess each step of the agent’s trajectory. Current self-evolution frameworks adopt these varying levels of granularity to tailor feedback mechanisms according to task complexity and the desired learning outcomes.
Outcome-based Reward Outcome-based Reward is a feedback mechanism that evaluates an agent based on the successful completion of predefined tasks. This reward is determined solely by the final state of the agent’s trajectory, regardless of the intermediate steps. A central challenge, particularly in dynamic environments like web or GUI navigation, is to effectively learn from both successful trajectories and the much more frequent failure trajectories. To address this, Direct Preference Optimization (DPO) rafailov2023direct is designed to directly maximize the likelihood of preferred responses while minimizing the KL-divergence with the reference policy. Similarly, RRHF yuan2023rrhf employs a ranking loss approach that aligns model probabilities of multiple responses with human preferences by ranking response probabilities without requiring auxiliary value models. Moreover, several works have developed specialized frameworks for agent self-evolution that are built upon outcome-based rewards. A straightforward approach is rejection sampling finetuning, as used in AutoWebGLM lai2024autowebglm . This method employs a pre-designed reward model to evaluate trajectory outcomes, identify the successful trajectories, and update the model with this high-quality data. DigiRL bai2024digirl models the GUI navigation task as a Markov Decision Process (MDP) and obtains a final, sparse reward at the end of an episode using a VLM-based evaluator. WebRL qi2024webrl develops a robust outcome-supervised reward model (ORM) to address the feedback sparsity inherent in dynamic web environments. The ORM evaluates task success within a self-evolving curriculum framework, enabling agents to learn from unsuccessful attempts and progressively improve.
Process-based Reward In contrast to outcome-based rewards, which provide a single, delayed signal, the process-based reward paradigm offers more precise and granular feedback by evaluating each step in an agent’s trajectory. Process-supervised reward models (PRMs) have been demonstrated to be significantly more reliable than outcome-supervised reward models (ORMs), particularly in domains requiring complex reasoning like solving math problems lightman2023let . However, obtaining such fine-grained step-level feedback traditionally requires extensive human annotations, which are both time-consuming and expensive to scale. To address this annotation bottleneck, Math-Shepherd wang2023math proposes an automatic process annotation framework that utilizes Monte Carlo Tree Search (MCTS) to gather step-wise supervision by assessing each step’s potential to derive the correct final answer. Similarly, AlphaMath chen2024alphamath trains a value model to evaluate the step correctness in solution paths and updates both the policy and value model through exploration and exploitation within an MCTS framework. By leveraging process-based rewards, agents can improve their capabilities in a progressive, step-by-step manner. rStar-Math guan2025rstar and AgentPRM choudhury2025process both propose methods to iteratively evolve the policy and the process reward model, generating progressively higher-quality reasoning paths without manual labels. Agent Q putta2024agentq integrates a step-wise verification mechanism into its MCTS process to collect high-quality trajectories, which are then used to iteratively refine the policy via DPO training.
Hybrid Reward The hybrid methods aim to provide more comprehensive learning signals by incorporating both the clarity of final task success (outcome-based) and the granular guidance of intermediate steps (process-based). These methods overcome the sparsity of outcome-only signals while grounding the agent’s step-by-step reasoning in the ultimate task goal. For example, GiGPO feng2025group addresses the instability of training long-horizon agents by introducing a dual-level reward mechanism. It provides an episode-level reward based on the final success of entire trajectories, while simultaneously assigning a localized, step-level reward for intermediate actions. This dual signal provides both a high-level directional goal and low-level corrective guidance. Similarly, SPA-RL wang2025spa proposes a reward decomposition method that bridges the gap between sparse outcome signals and dense process feedback. It attributes incremental progress to each step within multi-step trajectories based on the final task completion, effectively distributing the outcome-based reward across the process steps. This approach creates dense intermediate progress rewards that enhance reinforcement learning effectiveness while maintaining alignment with the ultimate task objectives.
5.5 Other Dimensions of Self-Evolution Methods
In addition to the core axes of learning paradigm, policy consistency, and reward granularity, Table 4 highlights several other important dimensions that differentiate self-evolution methods:
Feedback Type. The nature of feedback varies widely: reward-based methods leverage scalar rewards, natural language signals, or model confidence; imitation methods focus on demonstration trajectories and rationales; population-based methods use fitness scores or competitive signals. The feedback type fundamentally determines what information the agent uses to improve.
Data Source. Reward-based methods typically generate data through agent-environment interaction or engineered rules, while imitation learning often relies on human or expert-generated demonstrations. Population-based approaches draw from the collective experience of multiple agents or generations, enabling diverse exploration but requiring significant coordination.
Sample Efficiency. Imitation learning is generally the most sample-efficient, provided high-quality demonstrations are available, as agents can directly mimic expert behavior. Reward-based methods are moderately efficient, with efficiency highly sensitive to reward sparsity. Population-based evolution tends to be sample-inefficient, as it often requires evaluating a large number of agent variants through many trials.
Stability. Reward-based learning is sensitive to the quality and design of reward functions, risking reward hacking or unintended behaviors. Imitation learning depends heavily on the quality and diversity of demonstrations. Population-based methods are sensitive to population size and diversity, with small or homogeneous populations at risk of premature convergence.
Scalability. Scalability is determined by the feasibility of data or feedback collection and the ability to parallelize learning. Reward-based methods scale well when feedback is automated (e.g., via simulators). Imitation learning is often bottlenecked by the cost of collecting demonstrations. Population-based approaches can scale to large compute but are highly resource-intensive.
Together, these dimensions offer a more nuanced, multidimensional view of self-evolution strategies, guiding practitioners in selecting and designing agent learning pipelines that are best matched to the challenges of their specific domains.
6 Where to Evolve?
Self-evolving agents have facilitated advancements across a diverse array of domains and applications. Broadly, most of these applications can be systematically categorized into two groups: (1) general domain evolution, where agent systems evolve to expand their capabilities across a wide variety of tasks, mostly within the digital realm, and (2) specialized domain evolution, which evolves specifically to enhance their proficiency within particular task domains. In essence, evolution in general-purpose assistants focuses on transferring learned experience to a broader set of tasks, while evolution in specialized agents emphasizes deepening expertise within a specific domain.
6.1 General Domain Evolution
The first category, general domain evolution, refers to self-evolving agents designed for general-purpose applications, particularly as versatile digital assistants. These agents progressively enhance their capabilities to address a broad spectrum of user queries, especially in dynamic and diverse digital environments. Technically speaking, these general assistant agent enhance their abilities primarily via three mechanisms: memory optimization, curriculum-driven training, and model-agent co-evolution. These mechanisms collectively enable the agents to continuously adapt and effectively respond to increasingly complex user demands.
Memory Mechanism.
The most common mechanism facilitating agent evolution is the memory mechanism, wherein agents summarize historical success/failure experiences (wang2023voyager, ; zhang2024offline, ) into memory representations (zhang2024survey, ), anticipating that these distilled experiences will be beneficial when addressing previously unseen tasks. For instance, Mobile-Agent-E (wang2025mobile, ) employs a long-term memory structure consisting of "Tips," which provide general guidelines, and "Shortcuts," representing reusable action sequences derived from past experiences. This self-evolutionary module supports the continuous enhancement of performance on complex smartphone tasks. Another typical example is MobileSteward (liu2025mobilesteward, ), which coordinates multiple app-specific Agents under a central Agent, with specialized modules for task scheduling, execution, and evaluation. It also incorporates a memory-based self-evolution mechanism that summarizes successful executions to improve future cross-app instruction handling. Meanwhile, Generative Agents (park2023generative, ) store episodic memories of their experiences, synthesize higher-level reflections, and condition future planning on this self-reflection. In these examples, memory serves as the foundation that enables agents to internalize past experiences, abstract high-level patterns, and refine their future behavior.
Figure 8: Categorization of where to evolve into two major types: General Domain Evolution, which focuses on broad capability enhancement across diverse tasks (e.g., memory mechanisms, co-evolution, curriculum training), and Specific Domain Evolution, which targets domain-specific expertise in areas such as coding, GUI, finance, medical, education, and others.
Model-Agent Co-Evolution.
Another line of work is to perform Model-Agent Co-evolution for LLM agents. UI-Genie (xiao2025ui, ) constructs a specialized image-text reward model that scores trajectories at both step and task levels. It jointly fine-tunes the agent and reward model using synthetic trajectories—generated by controlled corruption and hard-negative mining—across multiple generations. WebEvolver (fang2025webevolver, ) introduces a co-evolving world model LLM that simulates web environments. It generates synthetic training data by predicting next observations and enables look-ahead reasoning during inference, which greatly improves real-web task success. Absolute Zero (zhao2025absolute, ) co-evolves a reasoning agent and its internal self-reward model through reinforced self-play. By adversarially generating increasingly challenging reasoning problems and optimizing the agent using internal self-certainty as a reward signal, the framework simultaneously updates both the agent’s policy and the self-rewarding mechanism. Together, these methods demonstrate the effectiveness of co-evolving agents and auxiliary models (e.g., reward or world models) to achieve more robust, generalizable, and scalable learning in LLM agentic systems.
Curriculum-Driven Training.
Curriculum-driven training also serves as a critical mechanism for building a self-evolving general assistant. For example, WebRL (qi2024webrl, ) uses a self-evolving curriculum: when an agent fails, similar but manageable tasks are automatically generated. Coupled with a learned reward model and adaptive policy updates, this yields a success rate uplift on WebArena benchmarks. Voyager (wang2023voyager, ) similarly leverages an automatic, bottom-up curriculum in Minecraft, where GPT-4 proposes appropriate next tasks based on agent progress, building a growing code-based skill library through iterative prompting and environmental feedback. These approaches highlight how curriculum learning enables agents to autonomously expand their capabilities through iterative task adaptation.
6.2 Specialized Domain Evolution
In addition to general digital agents, self-evolving agents have also been effectively applied within specialized domains, where their evolution is tailored to significantly enhance performance within narrower task sets.
Coding.
The power of self-evolving agents extends directly to practical applications like coding, where their ability to autonomously adapt and improve offers a transformative approach to software development. SICA (robeyns2025self, ) demonstrates that a self-improving coding agent can autonomously edit its own codebase and improve its performance on benchmark tasks. EvoMAC (hu2024self, ) introduces a self-evolving paradigm on multi-agent collaboration networks, which automatically optimizes individual agent prompts and multi-agent workflows, significantly improving code generation performance by overcoming the limitations of manually designed systems. AgentCoder (huang2024agentcoder, ) also focuses on a multi-agent code generation framework that self-evolves through iterative refinement. A programmer agent continuously improves code based on feedback from a test executor agent, validated against independent test cases from a test designer, significantly boosting effectiveness and efficiency. Zhang et al. (zhang2025adaptive, ) enable LLM agents to continuously evolve by filtering high-quality answers, stratifying earned experiences by difficulty, and adaptively selecting demonstrations from self-generated data, leading to significant performance improvements and the construction of ML libraries. While these instances differ in their specific mechanisms—ranging from single-agent self-editing to complex multi-agent collaborative networks and experience-based learning—they commonly share the core principle of iterative self-improvement and autonomous adaptation to enhance coding capabilities. These advancements highlight how self-evolving agents can dramatically enhance coding efficiency and code quality by continuously learning and optimizing.
Graphical User Interfaces (GUI).
Self-evolving GUI agents extend LLM capabilities from pure text reasoning to direct manipulation of desktop, web, and mobile interfaces, where they must cope with large discrete action spaces, heterogeneous layouts, and partial visual observability. Yuan et al. couple pixel-level vision with self-reinforcement, enabling the agent to iteratively refine click–type grounding accuracy without additional human labels (yuan2025enhancing, ). On real desktop software, the Navi agent from WindowsAgentArena replays and critiques its own failure trajectories, ultimately doubling its task-completion rate across 150 Windows challenges (bonatti2024windows, ). For open-web automation, WebVoyager fuses screenshot features with chain-of-thought reflection; successive self-fine-tuning raises its end-to-end success on unseen sites from 30 % to 59 % (he2024webvoyager, ), while ReAP adds episodic memories of past outcomes, recovering a further 29-percentage-point margin on previously failed queries (azam2025reap, ). Beyond RL and memory, AutoGUI continuously mines functionality annotations from live interfaces to expand a reusable skill library each training cycle (li2025autogui, ), and MobileUse deploys a hierarchical self-reflection stack that monitors, verifies, and revises smartphone actions in situ (wu2025mobileuse, ). Collectively, these systems epitomize the full triad of self-evolution— what evolves (grounding modules, skill memories), when it evolves (offline consolidation vs. online reflection), and how it evolves (reinforcement learning, synthetic data, hierarchical monitoring)—charting a path toward universally competent interface agents.
Financial.
The primary bottleneck in customizing agents for specialized domains like financial tasks lies in efficiently constructing and integrating a domain-specific knowledge base into the agent’s learning process—a challenge that can be effectively mitigated by incorporating self-evolving mechanisms. QuantAgent (wang2024quantagent, ) proposed a two-layer framework that iteratively refines the agent’s responses and automatically enhances its domain-specific knowledge base using feedback from simulated and real-world environments. This iterative process helps the agent progressively approximate optimal behavior, reduces reliance on costly human-curated datasets, and demonstrably improves its predictive accuracy and signal quality in trading tasks. TradingAgents (xiao2024tradingagents, ) incorporates dynamic processes such as reflection, reinforcement learning, and a feedback loop from real-world trading results, alongside collaborative debates, to continuously refine its strategies and enhance trading performance. These developments underscore the potential of self-evolving agents to revolutionize the financial domain by autonomously building domain expertise, adapting to dynamic market conditions, and continuously improving decision-making and trading performance.
Medical.
Self-evolving agents have become a powerful paradigm in medical AI, where adaptability and the ability to evolve are essential for managing the complexity and ever-changing nature of real-world clinical practice. One of the most prominent applications is hospital-scale simulation. For example, Agent Hospital (agent2024hospital, ) creates closed environments with LLM-driven doctors, patients, and nurses, allowing the doctor agent to treat thousands of virtual cases. This process helps these agents autonomously refine and evolve their diagnostic strategies without manual labeling, ultimately achieving strong performance on USMLE-style exams. Similarly, MedAgentSim (MedAgentSim2025, ) integrates an LLM doctor, patient, and tool agent. It records successful consultations as reusable trajectories and employs chain-of-thought reflection and consensus to drive self-evolution, improving success rates over successive interactions. Another example is EvoPatient (du2024evopatient, ) places a doctor agent and a patient agent in continuous dialogue. With each generation, they update their memory with high-quality exchanges: the patient develops more realistic symptom narratives, while the doctor learns to ask sharper questions. Notably, this happens without explicit gradient updates or hand-crafted rewards. Reinforcement learning is also central to building adaptive medical agents. For instance, DoctorAgent-RL (doctoragentRL2025, ) models consultations as a Markov decision process, using a reward function that scores diagnostic accuracy, coverage, and efficiency. This guides policy-gradient updates that help the agent ask more relevant questions and reach correct diagnoses faster than imitation-based approaches, thus achieving self-improvement. In addition, automated architecture-search approaches like Learning to Be a Doctor treat the workflow itself as an evolvable object, iteratively inserting specialist sub-agents or new reasoning hops to cover observed failure modes and improve multimodal diagnostic accuracy (zhuang2025archsearch, ). Finally, beyond clinical decision-making, self-evolving agents have also been extended to biomedical discovery. OriGene (zhang2025origene, ) functions as a virtual disease biologist that evolves by iteratively refining its analytical process. It leverages human and experimental feedback to update core reasoning templates, adjust tool usage strategies, and refine analytical protocols. Similarly, STELLA (jin2025stella, ) is a self-evolving biomedical research agent that improves over time by distilling successful reasoning workflows into reusable templates through its Template Library and expanding its Tool Ocean with external or newly assembled tools to meet emerging analytical needs.
Education.
Self-evolving LLM agents have also found strong applications in the education domain. At the learner level, self-evolving agents like the personalized tutor PACE (liu2025pace, ) adjust their prompts based on detailed student profiles and continually refine their questioning during conversations. Meanwhile, an LLM-to-LLM self-play framework generates diverse tutor–student dialogues that further fine-tune the agent, allowing its teaching strategies to evolve both during and after interactions. Another example is MathVC (yue2025mathvc, ), which employs symbolic persona profiles for virtual students and a meta-planner that orchestrates realistic problem-solving stages. This setup enables the agent’s conversational process to evolve step by step toward correct solutions, closely mirroring how collaborative learning naturally unfolds. On the instructor side, self-evolving agent systems like the professional-development platform i-vip (yang2025ivip, ) deploy a team of cooperating LLM agents—a coach, assessor, and feedback generator—that critique and enhance each other’s outputs in real time. These agents adapt their explanations based on teacher-learners’ responses and continue to evolve by incorporating expert feedback after deployment, thereby refining their prompt strategies over time Similarly, EduPlanner (zhang2025eduplanner, ) frames lesson-plan creation as an adversarial loop where a planner’s draft is repeatedly reviewed and refined by evaluator and optimizer agents until it meets diverse educational goals. Similarly, SEFL (zhang2025sefl, ) uses teacher–student self-play to generate large sets of homework–feedback examples, which then fine-tune a lightweight feedback model. This self-evolving process significantly improves the clarity and usefulness of the comments. Collectively, these examples illustrate how self-evolving LLM agents can dynamically adapt to both learners and instructors, driving more personalized, effective, and scalable educational experiences.
Others.
Beyond the four major verticals discussed above, self-evolving agents demonstrate broader applicability, delivering superior adaptability and performance in specialized domains where conventional agents often fall short. For instance, Arxiv Copilot (lin2024arxivcopilot, ) learns and adapts by incorporating historical user interactions, including generated answers, research trends, and ideas, into its thought database, enhancing its ability to provide personalized and augmented academic assistance. In a very different context, Voyager (wang2023voyager, ), an agent in the game Minecraft, excels at solving novel tasks from scratch in new worlds through a process of self-evolution. It continually refines its task goals via an automatic curriculum, expands its skill library, and enhances its actions using an iterative prompting mechanism without human intervention. Transitioning to domains that require explicit strategic planning, Agents-of-Change (belle2025agents, ) autonomously refines prompts and rewrites code based on iterative performance analysis and strategic research, thereby helping agents overcome inherent limitations in long-term strategic planning and achieve consistently superior and more coherent gameplay in complex environments like Settlers of Catan. Lastly, in the realm of diplomacy, Richelieu (zhao2024richelieu, ) introduces AI diplomacy agents that can self-evolve through their self-play mechanism, which allows the agent to augment its memory by acquiring diverse experiences without human data, thereby enhancing its strategic planning, reflection, and overall performance in diplomacy activities. While these diverse examples operate in distinct environments—from academic research and virtual game worlds to strategic board games and complex diplomatic negotiations—they all share the fundamental characteristic of leveraging continuous learning, self-refinement, and autonomous adaptation to achieve increasingly sophisticated and effective performance within their respective domains. These diverse examples reinforce the versatility of self-evolving agents, showcasing their growing potential to excel in a wide range of complex, dynamic, and human-like tasks beyond traditional domains.
7 Evaluation of Self-evolving Agents
Evaluating self-evolving agents presents a unique set of challenges that extend beyond the traditional assessment of static AI systems. Unlike conventional agents typically evaluated on a fixed set of tasks at a single point in time, self-evolving agents are designed to continuously learn, adapt, and improve through ongoing interaction with dynamic environments. Consequently, their evaluation must capture not only immediate task success but also crucial aspects such as adaptation over time, knowledge accumulation and retention, long-term generalization, and the ability to transfer learned skills across sequential or novel tasks, all while mitigating catastrophic forgetting. This demands a fundamental shift from conventional “single-shot” assessments to a longitudinal view of their growth trajectory.
Figure 9: Overview of evaluation angles for self-evolving agents, encompassing core Evaluation Goals and Metrics—such as adaptivity, retention, generalization, safety, and efficiency—and a continuum of Evaluation Paradigms spanning from static assessment to short-term adaptability and long-horizon lifelong learning evaluation.
7.1 Evaluation Goals and Metrics
To effectively evaluate self-evolving agents, we must move beyond traditional metrics and establish a comprehensive framework that captures their dynamic, adaptive, and long-term learning capabilities. A truly capable and desirable self-evolving agent must not only learn and improve but also remember past knowledge, transfer it to new situations, operate sustainably, and behave responsibly. Grounded in these critical requirements for continuous and robust AI, we categorize the key evaluation goals into five core dimensions: Adaptivity, Retention, Generalization, Efficiency, and Safety, as illustrated in Table 5. Each dimension addresses a vital aspect of an agent’s self-evolutionary process, providing a holistic view of its performance.
Adaptivity
Adaptivity serves as a foundational evaluation criterion for any self-evolving agent, measuring its ability to improve performance on in-domain tasks through experience. This dimension focuses on quantifying the learning curve and the extent of performance enhancement as an agent iterates and evolves within a specific domain. Rather than a static success rate, adaptivity is gauged over time, steps, or iterations. Typical metrics include the Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench , which tracks performance in downstream tasks as a function of the agent’s interaction history.
Retention
Retention is a crucial criterion for evaluating the stability of a self-evolving agent’s knowledge base. This dimension specifically focuses on the challenge of catastrophic forgetting, a common issue in lifelong learning where new knowledge acquisition erodes previously learned information, and knowledge retention within extended interactions. Two key metrics can be used to quantify this stability from different perspectives: Forgetting (FGT) and Backward Transfer (BWT) zheng2025lifelong . Specifically, Let $J_{i,t}$ be the performance of LLM agents on task $i$ after completing $t$ tasks. FGT and BWT can be calculated as follows:
$FGT_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}[\displaystyle\max_{j\in\{i,i+1,...,t\}}(\{J_{j,i}\}_{j})-J_{t,i}]\quad BWT_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}(J_{t,i}-J_{i,i})$
FGT evaluates the average accuracy drop on old tasks after an agent learns a new one, thereby measuring whether useful experience is successfully maintained. In contrast, BWT assesses the average accuracy improvement on old tasks due to the experience gained from a new task. A positive BWT indicates that new learning positively benefits old tasks, signifying successful knowledge transfer and a more robust, stable learning process.
Generalization
While Adaptivity and Retention focus on in-domain performance, Generalization is a pivotal measure of a self-evolving agent’s ability to apply its accumulated knowledge to new, unseen domains or tasks. A truly intelligent agent should not only perform well within its familiar territory but also demonstrate a capacity for cross-domain generalization. This capability can be evaluated by assessing an agent’s performance on a diverse set of tasks that span multiple task distributions and domains. Common approaches include computing aggregate performance metrics (e.g., mean success rates) across multi-domain test suites liu2023agentbench , sun2023adaplanner , and conducting out-of-domain evaluations using held-out task distributions that simulate real-world novelty scenarios hu2024agentgen ; peng2025dlpo .
Efficiency
Efficiency quantifies the resourcefulness of a self-evolving agent. As agents operate continuously and make decisions autonomously, it is essential to evaluate the cost and speed of their evolutionary process. These metrics are particularly important for practical, real-world applications where resources like computation and time are finite. Key indicators include token consumption hu2024treeplannerefficientcloselooptask , which measures the computational cost of an agent’s reasoning and generation steps), time consumption lu2024axis , the number of steps wang2023voyager and the number of interaction with the tools (e.g. tool productivity)  (wang2025otcpo, ) required to complete a task, which rewards agents for completing tasks in the fewest possible resource consumption. Several key metrics quantify agent efficiency in task execution, including token consumption (measuring computational overhead in reasoning and generation) hu2024treeplannerefficientcloselooptask , time expenditure lu2024axis , the number of required stepswang2023voyager , and tool interaction frequency (wang2025otcpo, ). These indicators collectively assess an agent’s ability to minimize resource utilization while maintaining task performance, with lower values generally reflecting more efficient operation.
Safety
From the perspective of self-evolving, the Safety domain critically examines whether these agents develop unsafe or undesirable behavioral patterns throughout their continuous evolution. This dimension assesses an agent’s adherence to predefined rules and its propensity for harmful actions. Key metrics in evaluating safety of self-evolving agents may include: (1) Safety Score zhang2024agentsafety , measures the proportion of test cases where the agent’s behavior is labeled “safe” ; (2) Harm Score andriushchenko2024agentharm , computes via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered; (3) Completion Under Policy (CuP) levy2024st , assesses whether an agent successfully completes a task while strictly adhering to a given set of rules or policies ; (4) Risk Ratio levy2024st , calculates the frequency of an agent’s rule violations along a specific dimension, providing a quantitative measure of non-compliance ; (5) Refusal Rate zhang2024agentasb ; andriushchenko2024agentharm , evaluates the proportion of tasks an agent refuses to perform due to their aggressive, malicious, or otherwise unsafe nature; (6) Leakage Rate shao2024privacylens , tracks how often an agent unintentionally leaks sensitive or private information.
Table 5: Overview of Agent Evaluation Metrics Across Core Dimensions
Goal Metric Description
Adaptivity Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench Performance in downstream tasks as a function of the agent’s interaction history
Adaptation Speed wang2023voyager How quickly an agent reaches a certain performance threshold or converges to an optimal strategy within a given adaptation period
Retention Forgetting (FGT) zheng2025lifelong The average accuracy drop on old tasks after an agent learns a new one, measuring whether useful experience is successfully maintained
Backward Transfer (BWT) zheng2025lifelong The average accuracy improvement on old tasks due to the experience gained from new tasks
Generalization Aggregate Performance liu2023agentbench ; sun2023adaplanner Mean success rates or other performance indicators across multi-domain test suites to gauge overall proficiency
Out-of-Domain (OOD) Performance hu2024agentgen ; peng2025dlpo The agent’s performance in held-out task distributions
Efficiency Token Consumption hu2024treeplannerefficientcloselooptask Computational overhead in reasoning and generation steps
Time Expenditure lu2024axis Total duration required for task completion
Number of Steps wang2023voyager Minimal actions needed to accomplish objectives
Tool Productivity wang2025otcpo The ratio between task benefit (e.g., answer accuracy) and tool usage cost (e.g., number of tool calls)
Safety Safety Score zhang2024agentsafety Proportion of test cases where agent behavior meets predefined safety criteria
Harm Score andriushchenko2024agentharm Graded assessment of harmful outputs based on violation severity
Completion Under Policy (CuP) levy2024st Task success rate while complying with specified constraints
Risk Ratio levy2024st Frequency of policy violations per interaction opportunity
Refusal Rate zhang2024agentasb ; andriushchenko2024agentharm Percentage of tasks declined due to safety concerns
Leakage Rate shao2024privacylens Incidence of unintended sensitive information disclosure
7.2 Evaluation Paradigm
The evaluation of self-evolving agents, given their continuous learning paradigm, necessitates a multi-faceted approach that extends beyond traditional static assessments. Current evaluation paradigm can be broadly categorized based on the temporal scope of the assessment: Static Assessment, Short-horizon Adaptive Assessment, and Long-horizon Lifelong Learning Ability Assessment. Each category addresses different aspects of an agent’s evolving capabilities, from its instantaneous performance to its long-term learning trajectory.
7.2.1 Static Assessment
Static assessment evaluates the instantaneous performance of self-evolving agents at a specific point in time. Although these agents are designed for continuous improvement, static methods remain crucial for establishing baseline performance, comparing different agent architectures on fixed task sets, or evaluating capabilities after discrete training phases. This approach aligns with conventional AI evaluation, focusing on immediate performance in fixed environments. While useful for assessing generalization in an “in-domain evolving, out-of-domain evaluation” paradigm, static assessment inherently does not capture the dynamic, continuous learning, or long-term evolutionary aspects central to self-evolving agents.
For evaluating an agent’s general capabilities at a given moment, standard benchmarks designed for static AI systems are often employed. These benchmarks offer diverse task domains and test various core agent competencies, providing a snapshot of an agent’s proficiency before or at specific stages of its evolution. These assessments can be systematically categorized into External Task-Solving Evaluation and Internal Agent Components Evaluation, where External Task-Solving Evaluation measures end-to-end performance in completing domain-specific or cross-domain tasks, and Internal Capability Evaluation focuses on fundamental components in the agent, including planning, tool utilization, memory management, multi-agent coordination, etc.
External Task-Solving Evaluation
This category assesses an agent’s end-to-end proficiency in completing tasks across various real-world or simulated environments. In scientific data analysis and machine learning engineering, benchmarks like ScienceAgentBenchchen2024scienceagentbench and MLE-Benchchan2024mle test agents’ ability to generate and execute code for data analysis and solve Kaggle-style problems. For web search/Browse, environments such as WebShopyao2022webshop , WebArenazhou2023webarena , X-WebAgentBench wang2025x , Mind2Webdeng2023mind2web , and BrowseCompwei2025browsecomp simulate realistic web interactions, complex Browse scenarios, and task completion under security constraints. In software engineering, the SWE-bench seriesjimenez2023swe ; swebenchverified ; aleithan2024swe ; yang2024swe uses real GitHub issues to assess agents’ code repair capabilities. For computer-use interactions, OSWorldxie2024osworld offers a unified environment for open-ended tasks involving various desktop and web applications. Specialized domains like marketing also feature benchmarks such as xbenchchen2025xbench . Beyond specific domains, generalist agent benchmarks like AgentBenchliu2023agentbench , GAIAmialon2023gaia , and TheAgentCompanyxu2024theagentcompany evaluate broad problem-solving abilities across multiple knowledge domains and professional tasks, simulating real-world demands on general AI assistants.
Internal Agent Components Evaluation
Beyond end-to-end task completion, assessing an agent’s underlying core competencies is crucial. These benchmarks evaluate fundamental capabilities that contribute to an agent’s overall intelligence and self-evolutionary potential. As for Planning, Benchmarks such as PlanBenchvalmeekam2023planbench , Natural Planzheng2024natural , AutoPlanBenchstein2025automating , and ACPBenchkokel2025acpbench comprehensively evaluate an agent’s ability to understand dynamic environments, devise strategies, decompose complex problems, and execute reasoning in various planning domains. For Tool Usage, simple benchmarks like ToolAlpacatang2023toolalpaca and ToolBenchqin2023toolllm test basic selection and parameter mapping, while more complex ones like ToolSandboxlu2024toolsandbox , Seal-Toolswu2024seal , API-Bankli2023api , T-Evalchen2023t , $\tau$-Benchyao2024tau , AceBenchchen2025acebench ) simulate real-world scenarios involving multi-turn interactions, implicit state dependencies, and nested calls. Memory Management benchmarks such as LTMbenchmarkcastillo-bolado2024beyond , MemoryAgentBenchhu2025evaluating , and StoryBenchwan2025storybench evaluate the agent’s capacity to retain and utilize information across multi-turn interactions, dynamic scenarios, and long-range dependencies. For evaluating Multi-Agent Collaboration, benchmarks such as MultiAgentBenchzhu2025multiagentbench and SwarmBenchruan2025benchmarking assess coordination, communication, and emergent swarm intelligence in both collaborative and competitive settings.
Metrics for Static Assessment
Typical metrics for static assessment include accuracy, success rate, progress rate, completion rate, and various domain-specific performance indicators (e.g., CodeBertScore, Valid Execution Rate, Pass Rate, F1 score). These metrics provide a singular performance score for an isolated invocation or a fixed set of tasks.
Table 6: Differences between Short-horizon Adaptive Assessment and Long-horizon Lifelong Learning Ability Assessment
Dimension Short-horizon Adaptation Assessment Long-horizon Lifelong Learning Ability Assessment
Primary Focus Immediate learning and incremental improvement within consistent or slightly varying tasks Continuous knowledge accumulation and sustained performance across diverse, evolving tasks and environments.
Core Challenges Rapid adaptation to minor changes; Improving on similar, repeated tasks Mitigating catastrophic forgetting; Robust knowledge transfer; Maintaining efficiency/safety over time; handling true novelty and significant distribution shifts
Temporal Scope Small number of sequential tasks or iterations over a short period; Improvement on the same or similar task types. Large, potentially unbounded sequence of diverse, cross-domain tasks; Very long interaction periods requiring integration of new skills with old
Table 7: Representative Benchmarks for Evaluating Self-Evolving Agents
Benchmark Name Task Domain Goal Core Metrics Task Quantity Temporal Scope
ScienceAgentBench chen2024scienceagentbench Scientific Data Analysis Adaptivity, Efficiency Valid Execution Rate, Success Rate, CodeBERTScore, API Cost 102 Static
MLE-Bench chan2024mle ML-Engineering Adaptivity - 75 Static
DS-Bench jing2025dsbenchfardatascience Data Science Adaptivity Task Success Rate, Cost, Inference Time, Competition-level Accuracy 540 Static
SWE-bench jimenez2023swe Software Engineering Adaptivity Pass Rate 2,294 Static
OSWorld xie2024osworld Computer-Use / GUI Adaptivity Success Rate 369 Static
Mobile-Eval-E wang2025mobile Computer-Use / GUI Adaptivity, Efficiency Action Accuracy, Reflection Accuracy, Termination Error 25 Static, Short-horizon
WebShop yao2022webshop Web Search / Browse Adaptivity Success Rate 12,087 Static
WebArena zhou2023webarena Web Search / Browse Adaptivity Success Rate 812 Static
WebWalkerQA wu2025webwalker Web Search / Browse Adaptivity, Efficiency Accuracy, Action Count 680 Static
ST-WebAgentBench levy2024st Web Search / Browse Safety Completion under Policy 235 Static
xbench chen2025xbench Web Search / Browse Adaptivity LLM-Judge Score 100 Static
BrowseComp wei2025browsecomp Web Search / Browse Adaptivity Accuracy 1,266 Static
Agent-SafetyBench zhang2024agentsafety General Safety Safety Sore 20,000 Static
LifelongAgentBench zheng2025lifelongagentbench General Adaptivity, Retention, Generalization Success Rate 1396 Long-horizon
AgentBench liu2023agentbench General Adaptivity, Generalization Success Rate, F1, Reward, Game Progress 1360 Static
GAIA mialon2023gaia General Adaptivity Accuracy 466 Static
TheAgentCompany xu2024theagentcompany General Adaptivity, Efficiency Completion Score, Steps, Cost paer Instance 175 Static
EvaLearn dou2025evalearnquantifyinglearningcapability General Adaptivity, Efficiency Accuracy, Slope, Position of 1st solution, Num of consecutive solutions 648 Long-Horizon
PlanBench valmeekam2023planbench Planning Adaptivity Accuracy $\sim$26,250 Static, Short-horizon
Natural Plan zheng2024natural Planning Adaptivity Exact Match 3,600 Static
ACPBench kokel2025acpbench Planning Adaptivity, Generalization Accuracy 3,720 Static
AppBench (wang-etal-2024-appbench, ) Planning Adaptivity Success Rate, F1 800 Static
ToolBench qin2023toolllm Tool Usage Adaptivity Pass Rate, Win Rate 126,486 Static
ToolSandbox lu2024toolsandbox Tool Usage Adaptivity Similarity Score 1,032 Static
Seal-Tools wu2024seal Tool Usage Adaptivity Accuracy, P/R/F1 14,076 Static
API-Bank li2023api Tool Usage Adaptivity Accuracy, Rouge 4,125 Static
T-Eval chen2023t Tool Usage Adaptivity Domain-Specific Score 23,305 Static
$\tau$-Bench yao2024tau Tool Usage Adaptivity Pass^k 165 Static
AceBench chen2025acebench Tool Usage Adaptivity Accuracy 2,000 Static
LTMBenchmark castillo-bolado2024beyond Agent Memory Retention, Efficiency Score, Accuracy, GoodAI LTM Score, Speed, Cost, Verbosity 30 Long-Horizon
StoryBench wan2025storybench Agent Memory Retention, Efficiency Accuracy, First-Try Accuracy, Longest Corr, Retry Count, Runtime Cost, Token Cons, 311 scene nodes, 86 choice nodes Short-Horizon, Long-Horizon
MemoryAgentBench hu2025evaluating Agent Memory Adaptivity SubEM, Recall, ROUGE F1, Accuracy, Recall@5, Model Based Acc/F1 2200 Static, Short-horizon
MultiAgentBench zhu2025multiagentbench Multi-Agent Collaboration Adaptivity KPI, Text-Based Score, Communication Score, Planing Score, Coordination Score 100 Static
SwarmBench ruan2025benchmarking Multi-Agent Collaboration Adaptivity Perspective-specific Metrics 5 Short-horizon
7.2.2 Short-Horizon Adaptive Assessment
Short-horizon adaptations extend beyond static evaluations by assessing an agent’s ability to adapt and improve over a relatively short period or a limited number of interactions. The agent might improve performance on the same task instance with more attempts, or adapt to new instances of the same task type. This category focuses on capturing the capacity of the self-evolving agent for immediate adaptability and incremental learning within a relatively consistent or slightly varying task distribution. These evaluation schemes can be broadly categorized into two ways: (1) augment traditional benchmarks with a temporal dimension, and (2) specially design benchmarks and metrics that can inherently support Short-Horizon dynamic learning.
Augmented Traditional Benchmarks
Many studies leverage existing benchmarks but introduce a new dimension to track performance over time. This typically involves analyzing performance as a function of the number of iterations, steps, or examples. For example, ADAShu2024automated evaluated the held-out test accuracy with the number of agent system iterations on the ARC benchmark chollet2019measure ; AWMwang2024agent studied the cumulative success rate over the process of online evaluation under WebArena map test splitzhou2023webarena , using a number of examples to mark the evolution progress; WebEvolverfang2025webevolver studied the success rate with self-improving iterations under Mind2web-Live pan2024webcanvas . This approach allows for tracking the Adaptivity of the agent within a confined scope.
Benchmarks with Built-in Dynamic Evaluation
Some benchmarks are designed with short-horizon dynamic learning in mind. MemoryAgentBench hu2025evaluating , for example, includes a “Test-Time Learning” (TTL) dimension that evaluates an agent’s ability to learn new tasks directly from conversation within a single interaction session. In practice, TTL is evaluated through two types of tasks: Multi-Class Classification and Recommendation. In these settings, the agent must utilize previously provided information—such as labeled examples in context or a long movie-related dialogue history—to perform new tasks like mapping sentences to class labels or recommending relevant movies. This assesses immediate adaptation and knowledge acquisition during ongoing interaction.
Metrics and Methods for Evaluating Short-Horizon Adaptations
The primary metrics and methods for short-horizon adaptations are designed to quantify Adaptivity. These include: (1) Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench , which tracks performance improvements as the agent interacts more with the environment or attempts a task multiple times. (2) Learning Curve Analysis, visualizing how performance (e.g., success rate, accuracy) changes over a limited number of training steps, episodes, or interactions hu2024automated ; wang2024agent . (3) Adaptation Speed wang2023voyager , measuring how quickly an agent reaches a certain performance threshold or converges to an optimal strategy within the short horizon.
Short-horizon adaptations are well-suited for evaluating the initial learning capabilities and immediate adaptability of self-evolving agents. They can effectively demonstrate whether an agent can learn from recent experiences and improve its performance on in-domain tasks. This category is widely used for current self-evolving agents. However, the limited temporal window makes it challenging to assess long-term knowledge retention (mitigating catastrophic forgetting) and true lifelong learning capabilities across vastly different or sequentially presented tasks.
7.2.3 Long-Horizon Lifelong Learning Ability Assessment
Long-horizon lifelong learning ability assessment is crucial for truly assessing self-evolving agents, as they focus on the agent’s ability to continuously acquire, retain, and reuse knowledge across diverse environments and over extended periods. As shown in Table 6, it mainly focuses on continuous learning, knowledge accumulation, and sustained performance across a diverse and potentially ever-changing stream of tasks or environments over an extended period. This is a nascent but critical area, where unique challenges include catastrophic forgetting, robust knowledge transfer across disparate tasks, efficient resource management over extended durations, and mitigating data leakage when continuously evaluating on evolving data distributions. Specialized benchmarks are emerging to tackle these complexities.
Currently, there are few benchmarks of this type. LTMBenchmark castillo-bolado2024beyond is a specialized benchmark focusing on long-term memory (LTM) evaluation. It assesses LLM agents’ memory retention and continual learning through dynamic conversational tests, using interleaved dialogues with controlled distractions to simulate real-world recall challenges. Key metrics include task accuracy, memory-span-weighted LTM Score, and efficiency measures (tests/hour, cost) for cross-architecture comparison. LifelongAgentBench zheng2025lifelongagentbench is another pioneering benchmark specifically designed to evaluate agent lifelong learning. It constructs sequences of interdependent tasks across domains like Database (DB), Operating System (OS), and Knowledge Graph (KG), requiring agents to progressively build upon previously acquired skills. This allows for systematic tracking of performance improvement and knowledge retention across a prolonged learning trajectory. In addition, there is a solution that constructs a dynamic benchmark through continuously updating benchmark datasets white2024livebench ; yang2025truly or evolving the benchmark itself by reconstructing original benchmarks to evaluate self-evolving agents, which can alleviate data leakage to some extent chen2025recent . Benchmark Self-Evolving wang2024benchmark , for example, proposes a solution to continuously update the existing benchmark through iteration. Preliminary findings from such dynamic benchmark scenarios have shown that model performance can degrade as the benchmark evolves, highlighting the difficulty of continuous adaptation.
Metrics for long-horizon lifelong learning go beyond simple success rates to quantify the agent’s evolving ability, such as Forgetting (FGT), Backward Transfer (BWT) zheng2025lifelong , Cost-per-Gain. Long-term Generalization metrics could involve assessing performance on a continuously evolving set of out-of-distribution tasks or measuring the breadth of tasks an agent can still perform effectively after prolonged learning across many domains.
Long-horizon lifelong learning ability assessment is essential for comprehensively evaluating the core promise of self-evolving agents: their ability to learn continuously, retain knowledge, and generalize effectively over extended periods. They are critical for assessing Retention, Generalization to truly novel scenarios, and the Efficiency of long-term operation. This area remains a key frontier for research in evaluating self-evolving agents.
8 Future Direction
8.1 Personalize AI Agents
With the increasing interest in self-evolving agents, deploying personalized agents has become a crucial and increasingly significant objective for the research community (zhang2024personalization, ). For instance, in applications such as chatbots, digital twins, and emotional support dialogues, a key challenge is enabling AI agents to accurately capture and adapt to users’ unique behavioral patterns or preferences over extended interactions. Existing personalized agents typically depend heavily on labeled data and post-training methodologies (cheng2024autopal, ). Recent work by zhang2025personalize proposes a self-generated preference data approach aimed at rapidly personalizing LLMs. TWIN-GPT wang2024twin leverages electronic health records to create digital twins of patients, enhancing the accuracy of clinical trial outcome predictions. However, these existing strategies hinge on the critical assumption that LLMs can consistently obtain high-quality, large-scale user data. In practical deployment scenarios, the primary challenge remains the cold-start problem: agents need to progressively refine their personalized understanding, accurately interpret user intentions, and effectively construct user profiles, even when initial data is limited. Additionally, significant challenges persist in personalized planning and execution, such as effective long-term memory management, external tool integration, and personalized generation (ensuring outputs consistently align with individual user facts and preferences) (li2025survey, ). Moreover, it is essential to ensure that self-evolving agents do not inadvertently reinforce or exacerbate existing biases and stereotypes, highlighting another critical direction for future research.
With the integration of personalized data, evaluation metrics for personalizing self-evolving agents should extend beyond intrinsic evaluations (e.g., directly assessing personalized generated text quality using metrics such as ROUGE (lin2004rouge, ) and BLEU (papineni2002bleu, )) or extrinsic evaluations (e.g., indirect assessments of personalization effects through recommendation systems, classification tasks, and other specific applications). Traditional personalization evaluation metrics often fail to adequately capture the evolving dynamics inherent in self-evolving agents. Consequently, future research calls for more lightweight and adaptive evaluation metrics (zhang2024personalization, ). Additionally, to better assess self-evolving personalized agents, there is a clear need for flexible, dynamic benchmarks capable of accurately evaluating agents’ performance, particularly in managing long-tailed personalization data throughout their self-evolving processes.
8.2 Generalization
Self-evolving agents also face considerable challenges in achieving robust generalization across diverse task domains and environments. The fundamental tension between specialization and broad adaptability remains one of the most pressing challenges in the field, with significant implications for scalability, knowledge transfer, and collaborative intelligence.
Scalable Architecture Design:
A central challenge in developing generalizable self-evolving agents lies in designing scalable architectures capable of maintaining performance as complexity and scope increase. Current agent systems frequently encounter a trade-off between specialization and generalization, where agents optimized for specific tasks struggle to transfer their learned behaviors to novel environments (chen2024optima, ). Additionally, the computational cost associated with dynamic reasoning in LLM-based agents grows non-linearly with the complexity of adaptation mechanisms, imposing practical constraints on achievable generalization within realistic resource limitations (kim2025cost, ). Recent studies indicate that self-evolving agents equipped with reflective and memory-augmented capabilities show substantial promise for enhancing generalization, particularly in smaller, resource-constrained models (liang2024sage, ). Nonetheless, these approaches continue to encounter limitations when addressing complex real-world scenarios that require sustained adaptation over prolonged periods.
Cross-Domain Adaptation:
Achieving generalization across domains represents a critical frontier for self-evolving agents. Current methods frequently rely on domain-specific fine-tuning, restricting agents’ adaptability to new environments without retraining (belle2025agents, ). Recent advancements in test-time scaling and inference-time adaptation provide promising pathways for enhancing cross-domain generalization (snell2024scaling, ; zhang2025survey, ). These techniques allow agents to dynamically allocate additional reasoning capacity to unfamiliar scenarios by scaling computational resources during inference, avoiding the need for increasing model parameters. Additionally, meta-learning strategies have demonstrated considerable potential in facilitating rapid few-shot adaptation to new domains (bilal2025meta, ). However, their effectiveness critically depends on an agent’s capability to accurately determine when supplementary computational resources are necessary and efficiently distribute these resources across diverse reasoning tasks.
Continual Learning and Catastrophic Forgetting:
Self-evolving agents must continuously adapt to new tasks while retaining previously acquired knowledge, a challenge exacerbated by the catastrophic forgetting phenomenon (ghosal2024understanding, ) of continual memorization (chen2024continual, ) inherent in LLMs (bell2025future, ). The stability-plasticity dilemma becomes particularly acute in foundation model-based agents, where the computational costs of retraining for every new task are prohibitive (zheng2025lifelong, ). Recent research has explored parameter-efficient fine-tuning methods, selective memory mechanisms, and incremental learning strategies to mitigate catastrophic forgetting while preserving adaptability (wang2024comprehensive, ). Nonetheless, achieving an optimal balance between efficiency and preventing model drift remains a significant open challenge, especially when agents operate under resource constraints or manage streaming data with stringent privacy considerations.
Knowledge Transferability:
Recent studies have identified critical limitations in knowledge transfer among AI agents. shi2025models emphasized that knowledge integration and transfer capabilities in current agents still require significant optimization. In particular, geng2025large found that LLM-based agents often fail to effectively propagate newly acquired knowledge from interactions to other agents, restricting their collaborative potential. Furthermore, vafa2025has revealed that foundation models might depend heavily on shallow pattern matching, rather than developing robust and transferable internal world models. These findings indicate several important future research directions: 1) it is essential to better understand the conditions under which knowledge acquired by one agent can be reliably generalized and communicated to others; 2) developing methods to quantify the limitations in agents’ knowledge transferability could lead to clearer insights into agent collaboration bottlenecks; 3) we need to have an explicit mechanism that encourage the formation of robust, generalizable world models could significantly improve the collaborative effectiveness of self-evolving agents.
8.3 Safe and Controllable Agents
As autonomous AI agents become increasingly capable of learning, evolving, and performing complex tasks independently, more agent-based studies are shifting their focus towards the deployment of safer and more controllable agents. These safety concerns arise primarily from user-related risks, such as vague or misleading instructions that lead agents to execute harmful actions, as well as environmental risks, including exposure to malicious content, such as phishing website links (zhou2025automating, ).
Many studies aimed to address safety concerns about the automatic adaptation of agents. For instance, TrustAgent (hua2024trustagent, ) implements pre-planning, in-planning, and post-planning strategies to foster safer agent behavior. However, as highlighted in (zharmagambetov2025agentdam, ), current agents based on LLM still struggle to accurately differentiate between sensitive information that is necessary and irrelevant information. A major challenge here is the precise identification and understanding of task-related versus unrelated information. Furthermore, managing agent actions when goals involve deceptive or unethical methods presents further difficulties, as ongoing learning uncertainty exacerbates these safety challenges for the deployment of controllable agents (anwar2024foundational, ). This uncertainty is reflected similarly in ambiguous contexts (bagdasarian2024airgapagent, ) and poorly designed memory modules (wang2025unveiling, ). Therefore, deploying a reliable, controllable, and safe self-evolving system has become a critical issue. Future research should focus on collecting larger-scale, more diverse real-world scenario data to support comprehensive learning of safe behaviors. Further refining the Agent Constitution by developing clearer, more understandable rules and case libraries is essential. Furthermore, exploring safer training algorithms and thoroughly investigating the impacts of privacy-protection measures on agent efficiency are necessary steps toward achieving a more balanced and secure deployment of autonomous AI agents.
8.4 Ecosystems of Multi-Agents
Multi-agent self-evolving systems face several unique challenges that require further exploration.
Balancing Individual and Collective Reasoning:
Recent studies highlight the difficulty of balancing independent reasoning with effective group decision-making in multi-agent environments (chen2025mdteamgpt, ; sun2025collab, ). While collective discussions can significantly enhance diagnostic reasoning, agents often risk becoming overly reliant on group consensus, thereby diminishing their independent reasoning capabilities. To mitigate this issue, future research should explore dynamic mechanisms that adjust the relative weight of individual versus collective input. Such an approach would help prevent decision-making from being dominated by a single or a small subset of agents, ultimately promoting robust, balanced consensus-building and innovation. Additionally, developing explicit knowledge bases and standardized updating methodologies—leveraging agents’ successes and failures—could further improve the agents’ self-evolution abilities and strengthen their individual reasoning contributions within collaborative contexts.
Efficient Frameworks and Dynamic Evaluation:
Another crucial challenge lies in developing efficient algorithms and adaptive frameworks that allow agents to collaborate effectively while preserving their individual decision-making strengths. (hu2024self, ) introduced adaptive reward models and optimized dynamic network structures, which can significantly enhance cooperative self-improvement among agents. However, a major gap identified by (sun2025collab, ) is the absence of clear mechanisms for agents to dynamically manage and update their knowledge. Addressing this issue will require new frameworks that explicitly integrate continuous learning and adaptive collaboration mechanisms. Furthermore, existing benchmarks for multi-agent evaluation are predominantly static (zhu2025multiagentbench, ) and therefore fail to capture the long-term adaptability and continuous evolution of agent roles. Future benchmarks should incorporate dynamic assessment methods, reflecting ongoing adaptation, evolving interactions, and diverse contributions within multi-agent systems, thus providing more comprehensive evaluation metrics for self-evolving agents.
9 Conclusion
The emergence of self-evolving agents marks a paradigm shift in artificial intelligence, moving beyond static, monolithic models toward dynamic agentic systems capable of continual learning and adaptation. As language agents are increasingly deployed in open-ended, interactive environments, the ability to evolve, adapting reasoning processes, tools, and behaviors in response to new tasks, knowledge, and feedback, has become essential for building the next generation of agentic systems. In this survey, we provide the first comprehensive and systematic review of self-evolving agents, organized around three foundational questions: what aspects of an agent should evolve, when evolution should occur, and how to implement evolutionary processes effectively. Moreover, we discuss several methods for evaluating the progress of self-evolving agents in terms of metrics and benchmarks, followed by corresponding applications and future directions. Looking ahead, realizing the full potential of self-evolving agents will be critical in laying the groundwork for Artificial Super Intelligence (ASI). The evolution of these agents will require significant advancements in models, data, algorithms, and evaluation practices, and so on. Addressing issues such as catastrophic forgetting, human preference alignment during autonomous evolution, and the co-evolution of agents and environments will be key to unlocking agents that are not only adaptive but also trustworthy and aligned with human values. We hope this survey provides a foundational framework for researchers and practitioners to design, analyze, and advance the development and progress of self-evolving agents.